### v3

# D2ST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

Abstract
Adapting large pre-trained image models to few-shot action recognition has proven to be an effective and efficient strategy for learning robust feature extractors, which is essential for few-shot learning. Typical fine-tuning based adaptation paradigm is prone to overfitting in the few-shot learning scenarios and offers little modeling flexibility for learning temporal features in video data. In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter), which is a novel adapter tuning framework well-suited for few-shot action recognition due to lightweight design and low parameter-learning overhead. It is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. In particular, we devise the anisotropic Deformable Spatio-Temporal Attention module as the core component of D2ST-Adapter, which can be tailored with anisotropic sampling densities along spatial and temporal domains to learn spatial and temporal features specifically in corresponding pathways, allowing our D2ST-Adapter to encode features in a global view in 3D spatio-temporal space while maintaining a lightweight design. Extensive experiments with instantiations of our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods for few-shot action recognition. Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition.
Keywords:
Action recognition Few-shot learning Adapter tuning
footnotetext: Code is available at https://github.com/qizhongtan/D2ST-Adapter.
Figure 1: Comparison with previous adapter-based action recognition methods. Both AIM [51] and DUALPATH [26] pre-learn the spatial and temporal features separately by duplicating the MSA modules, and insert vanilla adapter into each duplicated module for specific feature adaptation. ST-Adapter [25] first attempts to design the specialized adapter for video data using 3D convolution. Unlike these models, our D2ST-Adapter is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. Furthermore, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component, which configures different sampling density in the spatial and temporal domains to model two pathways specifically and enables D2ST-Adapter to perform feature adaptation in a global view in 3D spatio-temporal space while maintaining a lightweight design.
1 Introduction
Few-shot action recognition aims to learn an action recognition model from a set of base classes of video data, which has the capability to recognize novel categories of actions using only a few support samples. To this end, learning an effective feature extractor that is generalizable across different classes is crucial to the task. A typical way [2, 54, 1, 44, 42] is to leverage pre-trained large vision models [13, 37, 29] for feature learning by task adaptation.
Most existing methods [28, 43, 42] seek to adapt large vision models to few-shot action recognition by fine-tuning either the entire or partial model. While such methods have achieved promising performance, there are two important limitations that hamper them from exploiting the full potential of the pre-trained models. First, fully fine-tuning the pre-trained models is prone to overfitting in the few-shot learning scenarios whilst partial fine-tuning acquires limited effectiveness due to limited capacity for adaptation. Second, most large vision models are trained on image data due to the difficulty of collecting a large scale of video data. Thus, these fine-tuning based methods require to learn the temporal features as a post-process after feature learning by the pre-trained backbone, typically by constructing an auxiliary module to model the temporal dynamics [36, 46, 44, 43], which has limited effectiveness compared to the learning manner throughout the whole feature encoding stage.
As a prominent parameter-efficient fine-tuning technique [14, 58, 17, 12], adapter tuning [14, 4, 41] embeds learnable adapters with a negligible amount of parameters into pre-trained models flexibly while keeping the model frozen, conducting model adaptation efficiently and effectively. Compared to the fine-tuning adaptation paradigm, adapter tuning is particularly better-suited to few-shot learning scenarios. A straightforward way to apply adapter tuning to image-to-video transfer learning is to employ the vanilla adapter for image data directly to adapt the pre-learned spatial features and temporal features separately, which is exactly adopted by AIM [51] and DUALPATH [26]. As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules either in cascaded (AIM) or parallel (DUALPATH) manners, and insert vanilla adapter in each duplicated module for specific feature adaptation. However, such methods inevitably incur heavyweight model architecture and massive computation overhead. ST-Adapter [25] first attempts to design the specialized adapter for video data by inserting a 3D convolution layer into the vanilla adapter for learning spatio-temporal features while maintaining a lightweight structure. Nevertheless, we investigate two potential limitations of ST-Adapter. First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in the low-data scenarios, such a joint learning paradigm of spatio-temporal features for video data is inferior to learning the spatial and temporal features in a disentangled way, such as SlowFast [8] or two-stream design [31, 40]. Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
In this work we design the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations. Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in  Figure 1, in which the spatial pathway is responsible for capturing the spatial appearance features while the temporal pathway focuses on learning the temporal dynamics. Thus, our model can encode the spatio-temporal features in a disentangled manner. Moreover, we devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA) module as an essential component to model both the spatial and temporal pathways. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space, which first samples a group of reference points in the spatio-temporal space and then learns the offset for each point to shift them to more informative regions in the 3D space. These shifted points are used as key and value pairs for aDSTA to perform sparse self-attention, which allows our D2ST-Adapter to perform feature adaptation in a global view while maintaining a lightweight design.
A novel design of our aDSTA is that we make the sampling density anisotropic along the spatial and temporal domains by specific configuration, which allows for tailoring specialized versions of aDSTA to model the spatial and temporal pathways separately. In particular, we tailor aDSTA-T with denser sampling along the temporal domain than the spatial domain for the temporal pathway since it focuses on capturing temporal features. In contrast, the tailored aDSTA-S for the spatial pathway samples denser reference points along the spatial domain. To conclude, we make the following contributions:
  • We propose D2ST-Adapter, a novel adapter tuning method for few-shot action recognition, which is designed in a dual-pathway architecture to adaptively encode the spatial and temporal features in a disentangled manner. It can adapt large vision models pre-trained on image data to video data efficiently and effectively in few-shot learning scenarios due to the lightweight design.
  • We devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, allowing our D2ST-Adapter to encode features in a global view while maintaining a lightweight design.
  • Extensive experiments on five benchmarks with instantiations of our method on pre-trained ResNet [13] and ViT [29], demonstrate the superiority of our method over other methods, particularly in challenging scenarios like SSv2 benchmark where the temporal dynamics are critical for action recognition.
2 Related Work
Few-shot image classification. Existing works for few-shot image classification can be categorized into three groups: augmentation-based, optimization-based, and metric-based. Augmentation-based methods mainly hallucinate [45, 11] and generate [20, 55] new training samples adversarially to prevent the model from over-fitting. Optimization-based algorithms [9, 16, 30] aim to train a model that generalizes well and can be quickly adapted to new tasks through a few gradient update steps. Metric-based approaches recognize novel samples through nearest neighbor classification using a specific distance metric, such as cosine similarity [39, 52], Euclidean distance [32, 6, 53], and learnable metrics [34]. Our work follows the metric-based category due to its simplicity and effectiveness, and we focus on the more challenging video setting.
Few-shot action recognition. Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task. The first one is to investigate the spatio-temporal modeling [59, 60, 54, 27, 46, 56, 36, 43, 50, 47]. For example, MTFAN [46] proposes a motion encoder to learn global motion patterns and injects them into each video representations using a motion modulator. STRM [36] enriches local patch features and global frame features for joint spatio-temporal modeling. MoLo [43] designs a motion autodecoder to explicitly extract motion dynamics in a unified network. Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique. Another direction focuses on designing effective metric learning strategy [2, 1, 28, 44, 46, 15, 24, 57]. OTAM [2] utilizes the DTW [23] algorithm to calculate video distances with strict temporal alignment. TRX [28] exhaustively enumerates all sub-sequences of support and query videos and matches them using an attention mechanism. HyRSM [44] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints. We evaluate our D2ST-Adapter using the above three classical matching strategies to demonstrate its effectiveness and robustness.
Adapter tuning. As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [14], and quickly draws attention in many other research areas [4, 41, 12, 35]. A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model. During training, only the parameters of inserted adapters are updated while the original model remains frozen, leading to efficient task adaptation. Recently, some works apply this method to adapt image models for action recognition. AIM [51] duplicates multi-head self-attention modules and plugs adapters after them to separately learn the spatial and temporal features in a cascaded manner. Similarly, DUALPATH [26] explicitly builds the two-stream architecture upon ViT and utilizes adapters to parallelly learn the spatial and temporal features. ST-Adapter [25] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability. Our D2ST-Adapter follows the basic framework of ST-Adapter, and meanwhile optimize the essential technical designs for few-shot action recognition.
3 Method
3.1 Overview
Problem Formulation of Few-shot Action Recognition. The aim of few-shot action recognition is to learn an action recognizer from a set of base classes, which is able to recognize novel classes using only a few labeled samples. The task typically follows episodic paradigm [39, 2, 44, 43], which is consistent with general application scenarios for few-shot learning. An episode includes a support set $\mathcal{S}$ consisting of $N$ classes and $K$ labeled samples for each class (referred to as the $N$-way $K$-shot task), as well as a query set $\mathcal{Q}$ that contains unlabeled samples to be classified. In each episode, we aim to classify every query into one of the $N$ classes with the guidance of the support set. Such episodic task setting is consistently adopted during all the training, validation, and test stages.
Adapter Tuning Framework. We design a plug-and-play and lightweight adapter, dubbed D2ST-Adapter, which can be integrated into most existing large vision models. Thus we can leverage the powerful feature encoding capability of pre-trained large vision models by efficient image-to-video model adaptation while incurring only a small amount of parameter-tuning overhead. It is particularly well-suited to few-shot learning scenarios.
Figure 2 (a) illustrates the overall adapter tuning framework of our method. Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature transformation for task adaptation while maintaining the shape of feature maps consistent before and after each adaptation. Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen, leading to efficient task adaptation. The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy [32, 2, 43].
Our D2ST-Adapter is designed independently from the backbone structure of the feature extractor, thus it can be readily integrated into most of classical pre-trained deep vision models. In particular, we instantiate the backbone of feature extractor with two dominant deep visual learning frameworks, namely ResNet [13] and ViT [7], respectively.
Figure 2: (a) Overall adapter tuning framework of our method. (b) DST-Adapter, a convolutional version of our model, which also follows dual-pathway architecture, whereas both pathways are constructed based on 3D convolution. (c) D2ST-Adapter, which is designed in a dual-pathway architecture and both pathways are modeled based on the proposed aDSTA illustrated in (d).
3.2 D2ST-Adapter
Dual-pathway Adapter Architecture for Spatio-Temporal Disentanglement. Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features. Most large vision models are typically pre-trained on image data, thus our D2ST-Adapter should be capable of capturing both the spatial features and temporal features. To this end, we design the D2ST-Adapter as a dual-pathway architecture shown in Figure 2 (c), in which the spatial pathway is responsible for capturing the spatial semantics while the temporal pathway focuses on learning the temporal dynamics. As a result, our model is able to encode the spatio-temporal features for video data in a disentangled manner.
As a common practice of typical Adapters [14, 25], our D2ST-Adapter adopts the bottleneck architecture for reducing the computational complexity. It first downsamples the feature map into a low-dimensional feature space, then the downsampled features are fed into the spatial and temporal pathways concurrently to perform disentangled feature adaptation. Finally, both the adapted spatial and temporal features are fused by simple element-wise addition and upsampled back to the initial size. Formally, given the feature maps $\mathbf{F}_{i}^{\text{in}}\in\mathbb{R}^{T\times H\times W\times C}$ obtained from the $i$-th stage of the pre-trained backbone, containing $C$ channels of feature maps with spatial size $H\times W$ for each of $T$ frames, the feature adaptation by D2ST-Adapter can be formulated as:
$\mathbf{F}_{i}^{\text{out}}=\text{GELU}\left(\mathcal{F}_{\text{S}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus\mathcal{F}_{\text{T}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\right)\cdot\mathbf{W}_{\text{up}}.$ (1)
Herein, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{C\times C^{\prime}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{C^{\prime}\times C}$ are the transformation matrices of two linear layers for downsampling and upsampling, respectively. $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ denote the disentangled feature adaptation by the spatial and temporal pathways respectively. Below we discuss two feasible modeling mechanisms for $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$, which leverages 3D convolution and our proposed Anisotropic Deformable Spatio-Temporal Attention, respectively.
Modeling with 3D Convolution. A straightforward way to model the disentangled spatio-temporal feature adaptation ($\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ in Equation 1) is to employ 3D convolutional network, which can be tailored by configuring the shape of convolutional kernels to focus on learning either the spatial or the temporal features. Specifically, we model both $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ using 3D depth-wise convolutional layers followed by a batch normalization layer and a GELU layer, as shown in Figure 2 (b). The only difference between the modeling of them is that $\mathcal{F}_{\text{S}}$ uses $1\times 3\times 3$ convolutional kernel for capturing spatial features while $\mathcal{F}_{\text{T}}$ is constructed with $3\times 1\times 1$ convolutional layer to learn the temporal dynamics. The resulting version is called DST-Adapter.
While it has been validated that leveraging 3D convolutions with different kernel shapes can learn the spatial and temporal features separately [8, 38], An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain. Constructing deep convolutional network by stacking convolutional layers can enlarge the receptive field whilst increasing the model size and incurring heavy computational burden, which is inconsistent with the lightweight design principle for adapters.
Modeling with Anisotropic Deformable Spatio-Temporal Attention. To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise anisotropic Deformable Spatio-Temporal Attention (aDSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space. As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling density in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions. As a result, aDSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
We conduct several adaptations from the deformable attention in 2D space to our aDSTA in spatio-temporal 3D space. First, we employ Dynamic Position Embedding (DPE) [21] implemented as a 3D depth-wise convolution layer to learn the semantic-conditioned spatio-temporal position information for each token and incorporate it into the semantic features by simple element-wise addition. Thus, the feature map $\mathbf{F}^{\prime}_{i}\in\mathbb{R}^{T\times H\times W\times C^{\prime}}$ fused with the position information at the $i$-th stage is obtained by:
$\mathbf{F}^{\prime}_{i}=\mathcal{F}_{\text{DPE}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}}).$ (2)
Our aDSTA learns the offsets in 3D space for the reference points by a 3D convolutional network consisting of a 3D depth-wise convolution layer and a $1\times 1\times 1$ 3D convolution layer as well as a GELU in between. After shifting the reference points according to the learned offsets, we derive the features for each shifted reference point by performing trilinear interpolation among neighboring tokens within a 3D volume space around the point. For instance, the features for the shifted point $\mathbf{p}$ located at $(p_{t},p_{h},p_{w})$ is calculated via the trilinear interpolation $\mathcal{F}_{\text{tri-int}}$ by:
$\mathcal{F}_{\text{tri-int}}(\mathbf{p})=\sum_{\mathbf{r}}g(p_{t},r_{t})\cdot g(p_{h},r_{h})\cdot g(p_{w},r_{w})\cdot\mathbf{F}^{\prime}_{i}(\mathbf{r}),$ (3)
where $g(a,b)=\max(0,1-|a-b|)$ defines an interpolating cubic volume space centered at $a$ and $\mathbf{r}=(r_{t},r_{h},r_{w})$ indexes all tokens in the whole spatio-temporal 3D space. All shifted reference points are used as the keys and values shared for sparse self-attention. For instance, a token in $\mathbf{F}^{\prime}_{i}$ located at $(u_{t},u_{h},u_{w})$ serves as a query and the output $\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}$ of aDSTA is calculated by:
$\begin{split}&\mathbf{q}=\mathbf{F}^{\prime}_{i,(u_{t},u_{h},u_{w})}\cdot\mathbf{W}_{q},\mathbf{K}=\mathbf{P}\cdot\mathbf{W}_{k},\mathbf{V}=\mathbf{P}\cdot\mathbf{W}_{v},\\ &\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}=\text{softmax}(\mathbf{q}\mathbf{K}/\sqrt{C^{\prime}})\mathbf{V}\cdot\mathbf{W}_{o},\end{split}$ (4)
where $\mathbf{P}\in\mathbb{R}^{M\times C^{\prime}}$ is feature matrix of $M$ shifted points and $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ are projection matrices for the query, key, value and output respectively.
Anisotropic Sampling Density. The learned offsets for reference points are restricted in a limited range for two reasons. First, it can prevent the potential training collapse that all reference points are shifted to the same point. Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region. As a result, more sampling of reference points typically leads to more powerful and fine-grained representations by the shifted points. The deformable attention samples reference points uniformly in the 2D feature space with the same sampling density along two spatial dimensions. In contrast, we configure anisotropic sampling density in the spatial and temporal domains. Consequently, our D2ST-Adapter can leverage aDSTA to model both spatial feature adaptation $\mathcal{F}_{\text{S}}$ in the spatial pathway and temporal feature adaptation $\mathcal{F}_{\text{T}}$ in the temporal pathway. Specifically, we define the ‘Sampling kernel’ for aDSTA as:
Definition 1
Sampling kernel: formulated as the sampling densities of reference points along the dimensions of time, height and width, respectively, denoted as $<n_{t},n_{s},n_{s}>$, s.t. $M=n_{t}\times n_{s}\times n_{s}$. $M$ is the total number of sampled reference points.
Intuitively, an aDSTA module whose sampling kernel has large $n_{t}$ and small $n_{s}$ are more capable of learning the temporal features than learning the spatial features. On the other hand, sampling more reference points in the spatial domain than the temporal domain ($n_{s}>n_{t}$) generally makes aDSTA focus on learning the spatial features. Thus, we can tailor aDSTA by configuring the sampling kernel to model $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ correspondingly. In particular, we tailor two versions of aDSTA:
  • aDSTA-S for modeling $\mathcal{F}_{\text{S}}$ in the spatial pathway, whose sampling kernel has relatively large $n_{s}$ and small $n_{t}$ ($n_{s}>n_{t}$).
  • aDSTA-T for modeling $\mathcal{F}_{\text{T}}$ in the temporal pathway, which samples denser reference points in temporal domain than in spatial domain ($n_{t}>n_{s}$).
In practice, the values of $n_{t}$ and $n_{s}$ are tuned as hyper-parameters. Such modeling is analogous to classical property of 3D convolutional network that configuring the shape of the convolutional kernel can steer the convolutional operation to focus on learning features in different domain.
3.3 End-to-End Adapter Tuning
The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy. It first calculates the frame-wise $L_{2}$ distance matrix between the input query video and each class prototype derived by simply averaging the corresponding support samples. Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction. Formally, the similarity between query $q$ and class prototype $c$ can be expressed as:
$s(\mathbf{F}_{q},\mathbf{F}_{c})=\mathbf{M}([\mathbf{F}_{q}^{1},\cdots,\mathbf{F}_{q}^{T}],[\mathbf{F}_{c}^{1},\cdots,\mathbf{F}_{c}^{T}]),$ (5)
where $\mathbf{F}_{q}^{i}$ and $\mathbf{F}_{c}^{j}$ denote the features of the $i$-th frame in query $q$ and the $j$-th frame in class prototype $c$ while $\mathbf{M}$ is the matching metric such as OTAM [2] and Bi-MHM [44]. The whole model can be optimized in an end-to-end manner using the cross-entropy loss.
4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments on five standard few-shot action recognition benchmarks, including SSv2-Full [10], SSv2-Small [10], Kinetics [3], HMDB51 [19], and UCF101 [33]. The temporal features are crucial to the action recognition in SSv2-Full and SSv2-Small while the actions in other datasets relies more on the spatial features than the dynamic features. Following the typical data split [59, 2, 43], For SSv2-Full, SSv2-Small and Kinetics, we select 64/24/12 classes from the datasets as the training/validation/test set, respectively. As for HMDB51 and UCF101, we adopt the same data split as [54, 43].
Pre-trained backbone. Most prior arts employs the ResNet-50 [13] pre-trained on ImageNet [5] as backbone to extract frame features. Recently, CLIP-FSAR [42] conduct experiments on CLIP-ViT-B [7] and get remarkable performance. To verify the effectiveness of our D2ST-Adapter, we instantiate our method with both pre-trained backbones.
Implementation details. We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B) as shown in Figure 2 (a). We tune the sampling kernel for aDSTA-S and aDSTA-T for each backbone and the details are provided in the supplementary material. For a fair comparison with previous methods [2, 44, 43], we uniformly sample 8 frames (i.e., $T=8$) as the input of a video. We employ several standard techniques for data augmentation such as random crop and color jitter in the training stage. Bi-MHM [44] is used as the matching metric in comparison with other methods. Besides, OTAM [2] and TRX [28] are additionally used in the ablation study. We train the model episodically with Adam [18] optimizer. Following the common practice [44, 43], in the test stage, we randomly select 10,000 episodes from the test set and report the average accuracy.
[…]
Method Backbone Adaptation SSv2-Full SSv2-Small
paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot
CMN [59] ResNet-50 Full fine-tuning 34.4 -- -- -- 43.8 33.4 42.5 46.5
CMN-J [60] ResNet-50 Full fine-tuning -- -- -- -- -- 36.2 44.6 48.8
OTAM [2] ResNet-50 Full fine-tuning 42.8 49.1 51.5 52.0 52.3 36.4 45.9 48.0
ITANet [56] ResNet-50 Full fine-tuning 49.2 55.5 59.1 61.0 62.3 39.8 49.4 53.7
TRX [28] ResNet-50 Full fine-tuning 42.0 53.1 57.6 61.1 64.6 36.0 51.9 59.1
TA2[22] ResNet-50 Full fine-tuning 47.6 -- -- -- 61.0 -- -- --
MTFAN [46] ResNet-50 Full fine-tuning 45.7 -- -- -- 60.4 -- -- --
STRM [36] ResNet-50 Full fine-tuning 43.1 53.3 59.1 61.7 68.1 37.1 49.2 55.3
HyRSM [44] ResNet-50 Full fine-tuning 54.3 62.2 65.1 67.9 69.0 40.6 52.3 56.1
Nguyen et al. [24] ResNet-50 Full fine-tuning 43.8 -- -- -- 61.1 -- -- --
Huang et al. [15] ResNet-50 Full fine-tuning 49.3 -- -- -- 66.7 38.9 -- 61.6
HCL [57] ResNet-50 Full fine-tuning 47.3 54.5 59.0 62.4 64.9 38.7 49.1 55.4
SloshNet [49] ResNet-50 Full fine-tuning 46.5 -- -- -- 68.3 -- -- --
GgHM [50] ResNet-50 Full fine-tuning 54.5 -- -- -- 69.2 -- -- --
MoLo [43] ResNet-50 Full fine-tuning 56.6 62.3 67.0 68.5 70.6 42.7 52.9 56.4
ST-Adapter [25] ResNet-50 Adapter tuning 52.2 59.9 64.1 67.1 68.7 41.9 51.5 55.7
D2ST-Adapter (Ours) ResNet-50 Adapter tuning 57.0 65.2 69.5 71.4 73.6 45.8 56.6 60.9
CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 61.9 64.9 68.1 70.9 72.1 54.5 58.6 61.8
AIM [51] CLIP-ViT-B Adapter tuning 63.7 72.0 75.2 77.3 79.2 52.8 64.0 67.5
DUALPATH [26] CLIP-ViT-B Adapter tuning 64.5 72.6 76.0 78.0 79.8 53.5 64.4 68.1
ST-Adapter [25] CLIP-ViT-B Adapter tuning 64.2 72.4 76.0 77.4 79.5 53.1 64.2 68.0
D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 66.7 75.3 78.3 80.1 81.9 55.0 65.8 69.3
Figure 3: Illustration of actions our method excels at (1-4) and performs poorly at (5-8), respectively. The performance is evaluated on SSv2-Small benchmark in 1-shot setting.
Figure 4: Performance and efficiency of different methods with ResNet-50 backbone on SSv2-Small benchmark in 1-shot setting. Circle size indicates the memory usage.
4.2 Comparison with State-of-the-Art Methods
Benchmarks sensitive to temporal features. Table 1 shows the comparative results of different methods on SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition. Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin with both backbones in all few-shot settings except ‘5-shot’ on SSv2-Small. These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition. In particular, our D2ST-Adapter outperforms all existing adapter tuning methods for video data, including AIM, DUALPATH, and ST-Adapter, substantially on all settings, which reveals the benefit of two core technical designs of D2ST-Adapter, namely the dual-pathway architecture for disentangled learning of spatial and temporal features and anisotropic Deformable Spatio-Temporal Attention (aDSTA) for learning features in a global view.
To gain more insight into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter.  Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing poorly on simple actions that can be recognized primarily based on spatial features. Such results are consistent with the previous results that our model excels at recognizing actions involving complex temporal dynamics.
[…]
Method Backbone Adaptation Kinetics HMDB51 UCF101
paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot 1-shot 3-shot 5-shot
CMN [59] ResNet-50 Full fine-tuning 57.3 67.5 72.5 74.7 76.0 -- -- -- -- -- --
CMN-J [60] ResNet-50 Full fine-tuning 60.5 70.0 75.6 77.3 78.9 -- -- -- -- -- --
OTAM [2] ResNet-50 Full fine-tuning 73.0 75.9 78.7 81.9 85.8 54.5 65.7 68.0 79.9 87.0 88.9
ITANet [56] ResNet-50 Full fine-tuning 73.6 -- -- -- 84.3 -- -- -- -- -- --
TRX [28] ResNet-50 Full fine-tuning 63.6 76.2 81.8 83.4 85.9 53.1 66.8 75.6 78.2 92.4 96.1
TA2[22] ResNet-50 Full fine-tuning 72.8 -- -- -- 85.8 59.7 -- 73.9 81.9 -- 95.1
MTFAN [46] ResNet-50 Full fine-tuning 74.6 -- -- -- 87.4 59.0 -- 74.6 84.8 -- 95.1
STRM [36] ResNet-50 Full fine-tuning 62.9 76.4 81.1 83.8 86.7 52.3 67.4 77.3 80.5 92.7 96.9
HyRSM [44] ResNet-50 Full fine-tuning 73.7 80.0 83.5 84.6 86.1 60.3 71.7 76.0 83.9 93.0 94.7
Nguyen et al. [24] ResNet-50 Full fine-tuning 74.3 -- -- -- 87.4 59.6 -- 76.9 84.9 -- 95.9
Huang et al. [15] ResNet-50 Full fine-tuning 73.3 -- -- -- 86.4 60.1 -- 77.0 71.4 -- 91.0
HCL [57] ResNet-50 Full fine-tuning 73.7 79.1 82.4 84.0 85.8 59.1 71.2 76.3 82.5 91.0 93.9
SloshNet [49] ResNet-50 Full fine-tuning 70.4 -- -- -- 87.0 59.4 -- 77.5 86.0 -- 97.1
GgHM [50] ResNet-50 Full fine-tuning 74.9 -- -- -- 87.4 61.2 -- 76.9 85.2 -- 96.3
MoLo [43] ResNet-50 Full fine-tuning 74.0 80.4 83.7 84.7 85.6 60.8 72.0 77.4 86.0 93.5 95.5
ST-Adapter [25] ResNet-50 Adapter tuning 73.0 79.9 82.8 84.6 85.1 60.3 71.4 74.7 84.6 92.9 94.5
D2ST-Adapter (Ours) ResNet-50 Adapter tuning 75.0 81.4 84.3 85.8 87.0 61.6 73.0 76.6 86.9 94.4 95.6
CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 89.7 92.9 94.2 94.8 95.0 75.8 84.1 87.7 96.6 98.4 99.0
AIM [51] CLIP-ViT-B Adapter tuning 88.4 92.7 94.2 94.6 95.3 74.2 83.8 86.9 95.4 98.2 98.5
DUALPATH [26] CLIP-ViT-B Adapter tuning 88.8 92.8 94.3 94.7 95.4 74.9 84.6 87.5 95.7 98.4 98.7
ST-Adapter [25] CLIP-ViT-B Adapter tuning 88.5 92.6 94.0 94.6 95.1 74.1 84.3 87.3 95.9 98.5 98.9
D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 89.3 93.1 94.4 95.2 95.5 77.1 86.4 88.2 96.4 98.7 99.1
Benchmarks relying on spatial features. Table 2 presents the experimental results on the benchmarks relying more on spatial features, including Kinetics, HMDB51 and UCF101. Our method still performs best on most of the settings, which manifest the effectiveness and robustness of our method in these scenarios. We also observe that the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets. It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
Comparison of efficiency. We compare the efficiency between our D2ST-Adapter and other methods in terms of tunable parameter size and memory usage in  Figure 4. It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation. Meanwhile, our method outperforms these full fine-tuning based methods substantially while enjoying much fewer tunable parameters, which validates the superiority of our method.
4.3 Ablation study
We conduct ablation study on SSv2-Full and Kinetics datasets, using ResNet-50 as the pre-trained backbone and Bi-MHM [44] as the matching metric.
Comparison between different Adapters. We first compare the performance of different Adapters in Table 3. As shown in Figure 1, Vanilla-Adapter [14] only contains the downsampling and upsampling layers with a GELU nonlinearity in between. As described in Section 3.2, DST-Adapter is the convolutional version of our D2ST-Adapter which follows the dual-pathway structure but models spatial and temporal pathways with 3D convolution. The performance of full fine-tuning is also provided for reference. We make following observations.
  • Effect of learning temporal features by adapters. ST-Adapter outperforms Vanilla-Adapter distinctly, which indicates that using 3D convolution to learn joint spatio-temporal features facilitates the feature adaptation.
  • Effect of disentangled encoding of the spatial and temporal features. The performance improvement from ST-Adapter to DST-Adapter demonstrates the effectiveness of dual-pathway adapter architecture, which allows for encoding the spatial and temporal features in a disentangled manner.
  • Effect of the proposed anisotropic Deformable Spatial-Temporal Attention (aDSTA). Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on the SSv2-Full dataset, which is more challenging due to the necessity of encoding temporal features.
Table 3: Comparison between different adapters as well as full fing-tuning.
Method SSv2-Full Kinetics
1-shot 5-shot 1-shot 5-shot
Full Fine-tuning 44.9 57.0 72.4 84.3
Vanilla-Adapter 45.3 57.6 72.6 84.8
ST-Adapter 52.2 68.7 73.0 85.1
DST-Adapter (Ours) 53.7 69.6 73.3 85.5
D2ST-Adapter (Ours) 57.0 73.6 75.0 87.0
Table 4: Ablation study on the configuration of sampling kernel of aDSTA.
Configuration SSv2-Full Kinetics
1-shot 5-shot 1-shot 5-shot
aDSTA-Uniform 55.8 72.2 73.9 86.0
aDSTA-S & aDSTA-T 57.0 73.6 75.0 87.0
Table 5: Evaluation of generalization of our D2ST-Adapter across different matching metrics on SSv2-Full and Kinetics. The state-of-the-art methods for each metric are involved into comparison.
Matching Metric Method SSv2-Full Kinetics
1-shot 5-shot 1-shot 5-shot
RFPL [47] 47.0 61.0 74.6 86.8
OTAM [2] MoLo [43] 55.0 69.6 73.8 85.1
D2ST-Adapter 56.0 72.8 74.7 86.8
RFPL [47] 44.6 64.6 66.2 87.3
TRX [28] MoLo [43] 45.6 66.1 64.8 86.3
D2ST-Adapter 47.7 68.6 66.4 87.6
HyRSM [44] 54.3 69.0 73.7 86.1
Bi-MHM [44] MoLo [43] 56.6 70.6 74.0 85.6
D2ST-Adapter 57.0 73.6 75.0 87.0
Effect of configuration of the sampling kernel of aDSTA. To investigate the effect of configuring specialized sampling kernels for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway respectively, we additionally configure an aDSTA-Uniform module, which performs sampling with uniform density in both the spatial and temporal domains. Then we compare the performance of our D2ST-Adapter using two different settings, including 1) ‘aDSTA-Uniform’, in which aDSTA-Uniform is used to model both the spatial and temporal pathways and 2) ‘aDSTA-S & aDSTA-T’, in which D2ST-Adapter uses aDSTA-S and aDSTA-T to model the spatial and temporal pathways respectively. As shown in Table 4, the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling kernels for the spatial and temporal pathways.
Table 6: Performance of different methods using smaller backbones (i.e., ResNet-18 and ResNet-34) on SSv2-Full dataset.
Method ResNet-18 ResNet-34
1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 2-shot 3-shot 4-shot 5-shot
OTAM [2] 39.4 45.0 46.6 47.4 49.0 40.6 45.2 48.0 48.9 49.2
TRX [28] 29.9 38.2 44.0 48.2 50.3 32.4 41.6 47.7 52.0 53.5
HyRSM [44] 46.6 54.7 58.7 60.7 61.1 50.0 57.5 61.9 63.3 64.8
MoLo [43] 50.0 57.2 61.6 63.6 64.6 54.1 61.1 65.9 67.3 67.8
ST-Adapter [25] 47.3 54.3 58.1 61.3 62.2 48.9 55.6 59.4 62.5 64.1
DST-Adapter (Ours) 50.0 56.9 61.7 63.4 65.3 52.1 59.3 63.4 65.8 67.5
D2ST-Adapter (Ours) 53.0 60.4 65.0 67.1 68.6 54.4 62.5 66.0 69.1 70.6
Generalization across different matching metrics. To evaluate the generalization of our D2ST-Adapter across different matching metrics, we conduct the experiments using three classical matching metrics on SSv2-Full and Kinetics, including OTAM [2], TRX [28], and Bi-MHM [44], and compare our model with the state-of-the-art methods corresponding to each metric. Table 5 presents the results, which show that our D2ST-Adapter consistently outperforms other methods, especially on the challenging SSv2-Full benchmark, validating the well generalization of the proposed D2ST-Adapter across different matching metrics.
Instantiation on ResNet-18 and ResNet-34. To investigate the performance of our D2ST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings. Table 6 show that our D2ST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method. While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to D2ST-Adapter, which manifests the effectiveness of the proposed anisotropic Deformable Spatio-Temporal Attention (aDSTA).
Figure 5: Visualization of shifted reference points in both pathways for two video samples. Circle size indicates the importance for each point calculated by aggregating the attention from all queries.
4.4 Visualization
We visualize the shifted reference points learned by our D2ST-Adapter to investigate the effectiveness of the proposed aDSTA qualitatively. Following Deformable Attention [48], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance. Then we visualize top-100 most important shifted reference points in each pathway for a video sample, as shown in  Figure 5. Note that we sample 8 frames for all videos, if a reference point is shifted to a temporal position between two sampled frames, we visualize it on the nearest frame approximately.
The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways. Another interesting observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamic features. This is consistent with the design of specialized tailored aDSTA modules for different pathways by configuring the sampling kernels accordingly.
5 Conclusion
We present D2ST-Adapter, which is a novel adapter tuning method for few-shot action recognition. It is designed in a dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features. Moreover, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component of D2ST-Adapter, which enables it to encode features in a global view while maintaining lightweight design. Extensive experiments demonstrate that the superiority of our model over state-of-the-art methods, especially in challenging scenarios involving complex temporal features.
Supplementary Material
Appendix 0.A Tuning of the Sampling Kernels of aDSTA
The sampling kernels of aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter can be tuned on a held-out validation set. Generally, aDSTA-S should sample denser reference points in the spatial domain while aDSTA-T samples more points in the temporal domain. For the instantiation of our model with ResNet, we configure the sampling kernel of aDSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size. In contrast, we only need to tune one configuration for sampling kernel when using ViT as the backbone since the feature map always has fixed size in different stages. Table 7 shows the tuned configurations of the sampling kernels of aDSTA-S and aDSTA-T with different backbones. Besides, we also provide the configurations of sampling kernel for aDSTA-Uniform module which is constructed to validate the effect of configuring the sampling kernel of aDSTA in Table 4.
Table 7: Tuned sampling kernels (in the form of $(T,H,W)$) for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter. Besides, the sampling kernel for aDSTA-Uniform used in Table 4 for ablation study is also provided.
Backbone Feature Map Sampling Kernel Sampling Kernel Sampling Kernel
in aDSTA-S in aDSTA-T in aDSTA-Uniform
ResNet-50 (8, 56, 56) (2, 8, 8) (8, 4, 4) (4, 4, 4)
ResNet-50 (8, 28, 28) (2, 4, 4) (8, 2, 2) (4, 4, 4)
ResNet-50 (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
ResNet-50 (8, 7, 7) (2, 2, 2) (8, 1, 1) (4, 4, 4)
CLIP-ViT-B (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
Appendix 0.B Effect of the Inserted Position of D2ST-Adapter
Theoretically, our D2ST-Adapter can be inserted into any position of the backbone flexibly. To investigate the effect of the inserted position of D2ST-Adapter on the performance of the model, we conduct experiments with four different ways of inserting D2ST-Adapters into the pre-trained CLIP-ViT-B backbone (comprising 12 learning stages) on SSv2-Small dataset: a) early-insertion, which inserts the D2ST-Adapter into each of first 6 stages (close to the input), b) late-insertion that inserts the D2ST-Adapter into each of last 6 stages (close to the output), c) skip-insertion, which inserts the adapter into the backbone every two stages and d) full-insertion that inserts the adapter into each learning stage.
As shown in Table 8, late-insertion, namely inserting the proposed D2ST-Adapters into the last 6 stages, yields better performance than early-insertion, which implies that adapter tuning is more effective for task adaptation in deeper layers than in the shallower layers. It is reasonable since deeper layers generally capture the high-level semantic features, which are more relevant to task adaptation. Besides, full-insertion achieves the best performance at the expense of slightly more tunable parameters and memory usage.
Table 8: Effect of the inserted position of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset. Skip means adding D2ST-Adapter every other stage.
Insertion Position Tunable Parameters (%) Memory Usage 1-shot 5-shot
Early-insertion 4.1% 15.5GB 48.4 64.6
Late-insertion 4.1% 15.2GB 54.2 68.5
Skip-insertion 4.1% 15.2GB 53.3 67.9
Full-insertion 7.9% 16.7GB 55.0 69.3
Table 9: Effect of the bottleneck ratio of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset.
Ratio Tunable Parameters (%) Memory Usage 1-shot 5-shot
0.0625 1.4% 15.2GB 53.4 68.1
0.125 3.2% 15.7GB 54.2 68.6
0.25 7.9% 16.7GB 55.0 69.3
0.5 20.2% 18.7GB 54.7 69.4
Appendix 0.C Effect of the Bottleneck Ratio of D2ST-Adapter
The tunable parameter size is mainly determined by the bottleneck ratio of D2ST-Adapter, defined as the ratio of downsampled channel numbers to the initial size. Thus, we can balance between the model efficiency in terms of tunable parameter size and model effectiveness in terms of classification accuracy by tuning the bottleneck ratio. As shown in Table 9, larger bottleneck ratios typically yield more performance improvement while introducing more tunable parameters, and we set the bottleneck ratio to 0.25 in all the experiments based on the results.
Figure 6: Given a selected query within the salient object, top-50 relevant shifted reference points in terms of attention weights in both pathways are visualized for three video samples. The red stars denote the selected query tokens while the circles denote relevant shifted reference points.
Appendix 0.D More Visualizations
To provide more insight into the anisotropic Deformable Spatio-Temporal Attention (aDSTA), we provide more visualization results in Figure 6. To be specific, we manually select a point within the salient object as the query token and visualize top-50 relevant shifted reference points to the query in terms of attention weights. The results show that our model is able to capture the salient object through the shifted reference points in both spatial and temporal domains.

### v4

@@
 Abstract
-Adapting large pre-trained image models to few-shot action recognition has proven to be an effective and efficient strategy for learning robust feature extractors, which is essential for few-shot learning.
-Typical fine-tuning based adaptation paradigm is prone to overfitting in the few-shot learning scenarios and offers little modeling flexibility for learning temporal features in video data.
-In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter), which is a novel adapter tuning framework well-suited for few-shot action recognition due to lightweight design and low parameter-learning overhead.
-It is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner.
-In particular, we devise the anisotropic Deformable Spatio-Temporal Attention module as the core component of D2ST-Adapter, which can be tailored with anisotropic sampling densities along spatial and temporal domains to learn spatial and temporal features specifically in corresponding pathways, allowing our D2ST-Adapter to encode features in a global view in 3D spatio-temporal space while maintaining a lightweight design.
-Extensive experiments with instantiations of our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods for few-shot action recognition.
+Adapting pre-trained image models to video modality has proven to be an effective strategy for robust few-shot action recognition.
+In this work, we explore the potential of adapter tuning in image-to-video model adaptation and propose a novel video adapter tuning framework, called Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter).
+It features a lightweight design, low adaptation overhead and powerful spatio-temporal feature adaptation capabilities.
+D2ST-Adapter is structured with an internal dual-pathway architecture that enables built-in disentangled encoding of spatial and temporal features within the adapter, seamlessly integrating into the single-stream feature learning framework of pre-trained image models.
+In particular, we develop an efficient yet effective implementation of the D2ST-Adapter, incorporating the specially devised anisotropic Deformable Spatio-Temporal Attention as its pivotal operation.
+This mechanism can be individually tailored for two pathways with anisotropic sampling densities along the spatial and temporal domains in 3D spatio-temporal space, enabling disentangled encoding of spatial and temporal features while maintaining a lightweight design.
+Extensive experiments by instantiating our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods.
 Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition.
-Keywords:
-Action recognition Few-shot learning Adapter tuning
-footnotetext: Code is available at https://github.com/qizhongtan/D2ST-Adapter.
-Figure 1: Comparison with previous adapter-based action recognition methods.
-Both AIM [51] and DUALPATH [26] pre-learn the spatial and temporal features separately by duplicating the MSA modules, and insert vanilla adapter into each duplicated module for specific feature adaptation.
-ST-Adapter [25] first attempts to design the specialized adapter for video data using 3D convolution.
-Unlike these models, our D2ST-Adapter is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner.
-Furthermore, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component, which configures different sampling density in the spatial and temporal domains to model two pathways specifically and enables D2ST-Adapter to perform feature adaptation in a global view in 3D spatio-temporal space while maintaining a lightweight design.
+Code is available at https://github.com/qizhongtan/D2ST-Adapter.
+footnotetext: * Equal contribution.   $\dagger$ Corresponding author.
+Figure 1: Comparison with previous adapter-based methods.
+Both AIM and DUALPATH pre-learn the spatial and temporal features separately by duplicating the whole MSA modules, and insert vanilla adapter into each duplicated module for feature adaptation, which incurs heavy-weight model design.
+ST-Adapter designs the adapter for video data using 3D convolution whose performance is limited by the joint learning of spatio-temporal features.
+In contrast, our D2ST-Adapter is designed in a dual-pathway architecture to adapt the spatial and temporal features in a disentangled manner while maintaining the single-stream feature learning framework of the pre-trained model.
+Furthermore, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which configures anisotropic sampling densities in the spatial and temporal domains to model two pathways specifically and enables D2ST-Adapter to adapt features in a global view.
 1 Introduction
-Few-shot action recognition aims to learn an action recognition model from a set of base classes of video data, which has the capability to recognize novel categories of actions using only a few support samples.
-To this end, learning an effective feature extractor that is generalizable across different classes is crucial to the task.
-A typical way [2, 54, 1, 44, 42] is to leverage pre-trained large vision models [13, 37, 29] for feature learning by task adaptation.
+Few-shot action recognition aims to learn an action recognition model from a set of base classes of video data, which can recognize novel categories of actions using only a few support samples.
+To this end, learning an effective feature extractor that is generalizable across different classes is crucial.
+A typical way [3, 45, 1, 33, 35] is to leverage pre-trained large image models [10, 29, 24] for feature learning by image-to-video model adaptation.
 Most existing methods [23, 34, 35] seek to adapt large vision models to few-shot action recognition by fine-tuning either the entire or partial model.
 While such methods have achieved promising performance, there are two important limitations that hamper them from exploiting the full potential of the pre-trained models.
-First, fully fine-tuning the pre-trained models is prone to overfitting in the few-shot learning scenarios whilst partial fine-tuning acquires limited effectiveness due to limited capacity for adaptation.
-Second, most large vision models are trained on image data due to the difficulty of collecting a large scale of video data.
-Thus, these fine-tuning based methods require to learn the temporal features as a post-process after feature learning by the pre-trained backbone, typically by constructing an auxiliary module to model the temporal dynamics [36, 46, 44, 43], which has limited effectiveness compared to the learning manner throughout the whole feature encoding stage.
-As a prominent parameter-efficient fine-tuning technique [14, 58, 17, 12], adapter tuning [14, 4, 41] embeds learnable adapters with a negligible amount of parameters into pre-trained models flexibly while keeping the model frozen, conducting model adaptation efficiently and effectively.
-Compared to the fine-tuning adaptation paradigm, adapter tuning is particularly better-suited to few-shot learning scenarios.
-A straightforward way to apply adapter tuning to image-to-video transfer learning is to employ the vanilla adapter for image data directly to adapt the pre-learned spatial features and temporal features separately, which is exactly adopted by AIM [51] and DUALPATH [26].
-As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules either in cascaded (AIM) or parallel (DUALPATH) manners, and insert vanilla adapter in each duplicated module for specific feature adaptation.
-However, such methods inevitably incur heavyweight model architecture and massive computation overhead.
-ST-Adapter [25] first attempts to design the specialized adapter for video data by inserting a 3D convolution layer into the vanilla adapter for learning spatio-temporal features while maintaining a lightweight structure.
+First, fully fine-tuning pre-trained models tends to overfit in few-shot learning scenarios, whereas partial fine-tuning is less effective due to its limited capacity for adaptation.
+Second, most large vision models are trained on image data due to the scarcity of video data.
+Thus, these fine-tuning based methods require to learn the temporal features as a post-processing step after feature extraction by the pre-trained backbone, typically by constructing an auxiliary module to model the temporal dynamics [28, 38, 33, 34], which has limited effectiveness compared to methods that incorporate temporal learning throughout the entire feature learning stage.
+As a prominent parameter-efficient fine-tuning technique [11, 50, 13, 9], adapter tuning [11, 4, 32] is particularly well-suited to few-shot learning scenarios due to much lower parameter-learning overhead.
+A straightforward way to apply adapter tuning to image-to-video model adaptation is to employ the vanilla image adapter [11] to adapt the pre-learned spatial and temporal features separately, which is exactly adopted by AIM [44] and DUALPATH [22].
+As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules, either in cascaded manner (AIM) or parallel manner (DUALPATH).
+Subsequently, they insert individual vanilla adapters in each duplicated module for specific feature adaptation.
+However, such methods inevitably result in a heavyweight model architecture and substantial computational overhead.
+As illustrated in Figure 1, ST-Adapter [21] is the first to design a specialized adapter for video data by incorporating a 3D convolution layer into the vanilla adapter to learn spatio-temporal features.
 Nevertheless, we investigate two potential limitations of ST-Adapter.
-First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in the low-data scenarios, such a joint learning paradigm of spatio-temporal features for video data is inferior to learning the spatial and temporal features in a disentangled way, such as SlowFast [8] or two-stream design [31, 40].
+First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in low-data scenarios, such a joint learning paradigm for spatio-temporal features in video data is inferior to learning the spatial and temporal features in a disentangled manner, such as SlowFast [6] and other two-stream designs [25, 31].
 Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
-In this work we design the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations.
-Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in  Figure 1, in which the spatial pathway is responsible for capturing the spatial appearance features while the temporal pathway focuses on learning the temporal dynamics.
-Thus, our model can encode the spatio-temporal features in a disentangled manner.
-Moreover, we devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA) module as an essential component to model both the spatial and temporal pathways.
-It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space, which first samples a group of reference points in the spatio-temporal space and then learns the offset for each point to shift them to more informative regions in the 3D space.
-These shifted points are used as key and value pairs for aDSTA to perform sparse self-attention, which allows our D2ST-Adapter to perform feature adaptation in a global view while maintaining a lightweight design.
-A novel design of our aDSTA is that we make the sampling density anisotropic along the spatial and temporal domains by specific configuration, which allows for tailoring specialized versions of aDSTA to model the spatial and temporal pathways separately.
-In particular, we tailor aDSTA-T with denser sampling along the temporal domain than the spatial domain for the temporal pathway since it focuses on capturing temporal features.
-In contrast, the tailored aDSTA-S for the spatial pathway samples denser reference points along the spatial domain.
+In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations.
+Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in Figure 1, in which the spatial pathway is responsible for capturing the appearance features while the temporal pathway focuses on learning the temporal dynamics.
+This built-in disentangled encoding of spatio-temporal features equips our D2ST-Adapter with superior image-to-video model adaptation capabilities while maintaining a lightweight design and low adaptation overhead.
+As a result, it benefits from a more efficient model design compared to AIM and DUALPATH, which perform separate feature pre-learning and individual adaptation for spatial and temporal features by duplicating learning modules.
+Meanwhile, our D2ST-Adapter demonstrates superior feature learning capability for video data compared to ST-Adapter, owing to its disentangled adaptation scheme.
+We develop an efficient yet effective implementation of the D2ST-Adapter, featuring the specially devised anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the pivotal operation to enable built-in disentangled encoding capability.
+It adapts the deformable attention [40] from 2D image space to 3D spatio-temporal space, which allows our D2ST-Adapter to encode features in both spatial and temporal domains in a global view while keeping lightweight modeling.
+One novel design of our aDSTA is that it can be tailored with anisotropic sampling densities along spatial and temporal domains, enabling specialized versions of aDSTA to model the spatial and temporal pathways separately.
+As shown in Figure 1, we tailor aDSTA-T with denser sampling along temporal domain than spatial domain for temporal pathway since it focuses on capturing temporal features.
+In contrast, the tailored aDSTA-S for spatial pathway samples denser points along spatial domain.
 To conclude, we make the following contributions:
-  • We propose D2ST-Adapter, a novel adapter tuning method for few-shot action recognition, which is designed in a dual-pathway architecture to adaptively encode the spatial and temporal features in a disentangled manner.
-It can adapt large vision models pre-trained on image data to video data efficiently and effectively in few-shot learning scenarios due to the lightweight design.
-  • We devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, allowing our D2ST-Adapter to encode features in a global view while maintaining a lightweight design.
+  • We propose D2ST-Adapter, a novel video adapter tuning framework for few-shot action recognition.
+It is designed in a dual-pathway architecture featuring built-in disentangled adaptation for spatial and temporal features, enabling effective yet efficient image-to-video model adaptation for few shot action recognition.
+  • We devise aDSTA, which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, yielding a lightweight implementation of D2ST-Adapter.
   • Extensive experiments on five benchmarks with instantiations of our method on pre-trained ResNet [10] and ViT [24], demonstrate the superiority of our method over other methods, particularly in challenging scenarios like SSv2 benchmark where the temporal dynamics are critical for action recognition.
 2 Related Work
-Few-shot image classification.
-Existing works for few-shot image classification can be categorized into three groups: augmentation-based, optimization-based, and metric-based.
-Augmentation-based methods mainly hallucinate [45, 11] and generate [20, 55] new training samples adversarially to prevent the model from over-fitting.
-Optimization-based algorithms [9, 16, 30] aim to train a model that generalizes well and can be quickly adapted to new tasks through a few gradient update steps.
-Metric-based approaches recognize novel samples through nearest neighbor classification using a specific distance metric, such as cosine similarity [39, 52], Euclidean distance [32, 6, 53], and learnable metrics [34].
-Our work follows the metric-based category due to its simplicity and effectiveness, and we focus on the more challenging video setting.
 Few-shot Action Recognition.
 Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task.
-The first one is to investigate the spatio-temporal modeling [59, 60, 54, 27, 46, 56, 36, 43, 50, 47].
-For example, MTFAN [46] proposes a motion encoder to learn global motion patterns and injects them into each video representations using a motion modulator.
+The first one is to investigate the spatio-temporal modeling [52, 45, 28, 34, 43, 39, 36, 16].
 STRM [28] enriches local patch features and global frame features for joint spatio-temporal modeling.
@@
 Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique.
-Another direction focuses on designing effective metric learning strategy [2, 1, 28, 44, 46, 15, 24, 57].
+Another direction focuses on designing effective metric learning strategies [3, 23, 33, 38, 12, 20].
 OTAM [3] utilizes the DTW [19] algorithm to calculate video distances with strict temporal alignment.
-TRX [28] exhaustively enumerates all sub-sequences of support and query videos and matches them using an attention mechanism.
+TRX [23] exhaustively enumerates all sub-sequences of support and query videos and matches them using attention mechanism.
 HyRSM [33] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints.
-We evaluate our D2ST-Adapter using the above three classical matching strategies to demonstrate its effectiveness and robustness.
+We evaluate our D2ST-Adapter using the above three classical matching metrics to demonstrate its effectiveness and robustness.
 Adapter Tuning.
-As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [14], and quickly draws attention in many other research areas [4, 41, 12, 35].
+As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [11], and quickly draws attention in many other research areas [4, 32, 9].
 A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model.
@@
 AIM [44] duplicates multi-head self-attention modules and plugs adapters after them to separately learn the spatial and temporal features in a cascaded manner.
-Similarly, DUALPATH [26] explicitly builds the two-stream architecture upon ViT and utilizes adapters to parallelly learn the spatial and temporal features.
+Similarly, DUALPATH [22] explicitly builds two-stream architecture upon ViT and utilizes adapters to parallelly learn spatial and temporal features.
 ST-Adapter [21] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability.
@@
 Problem Formulation of Few-shot Action Recognition.
-The aim of few-shot action recognition is to learn an action recognizer from a set of base classes, which is able to recognize novel classes using only a few labeled samples.
-The task typically follows episodic paradigm [39, 2, 44, 43], which is consistent with general application scenarios for few-shot learning.
+The task of few-shot action recognition typically follows episodic paradigm [30, 3, 33, 34].
 An episode includes a support set $\mathcal{S}$ consisting of $N$ classes and $K$ labeled samples for each class (referred to as the $N$-way $K$-shot task), as well as a query set $\mathcal{Q}$ that contains unlabeled samples to be classified.
@@
 Adapter Tuning Framework.
-We design a plug-and-play and lightweight adapter, dubbed D2ST-Adapter, which can be integrated into most existing large vision models.
-Thus we can leverage the powerful feature encoding capability of pre-trained large vision models by efficient image-to-video model adaptation while incurring only a small amount of parameter-tuning overhead.
-It is particularly well-suited to few-shot learning scenarios.
+We design a plug-and-play and lightweight adapter for video data, dubbed D2ST-Adapter, which can be integrated into most existing large image models.
+Thus we can leverage the powerful feature encoding capability of pre-trained models by efficient image-to-video model adaptation with only a small amount of parameter-tuning overhead.
+It is particularly well-suited to few-shot scenarios.
 Figure 2 (a) illustrates the overall adapter tuning framework of our method.
-Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature transformation for task adaptation while maintaining the shape of feature maps consistent before and after each adaptation.
-Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen, leading to efficient task adaptation.
-The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy [32, 2, 43].
-Our D2ST-Adapter is designed independently from the backbone structure of the feature extractor, thus it can be readily integrated into most of classical pre-trained deep vision models.
-In particular, we instantiate the backbone of feature extractor with two dominant deep visual learning frameworks, namely ResNet [13] and ViT [7], respectively.
+Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature adaptation.
+Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen.
+The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategies [3, 23, 33].
+We instantiate the backbone of feature extractor with ResNet and ViT respectively.
 Figure 2: (a) Overall adapter tuning framework of our method.
-(b) DST-Adapter, a convolutional version of our model, which also follows dual-pathway architecture, whereas both pathways are constructed based on 3D convolution.
+(b) DST-Adapter, a convolutional version of our model, which follows dual-pathway architecture, whereas both pathways are constructed based on 3D convolution.
 (c) D2ST-Adapter, which is designed in a dual-pathway architecture and both pathways are modeled based on the proposed aDSTA illustrated in (d).
 3.2 D2ST-Adapter
-Dual-pathway Adapter Architecture for Spatio-Temporal Disentanglement.
+Dual-pathway Adapter Architecture.
 Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features.
@@
 Formally, given the feature maps $\mathbf{F}_{i}^{\text{in}}\in\mathbb{R}^{T\times H\times W\times C}$ obtained from the $i$-th stage of the pre-trained backbone, containing $C$ channels of feature maps with spatial size $H\times W$ for each of $T$ frames, the feature adaptation by D2ST-Adapter can be formulated as:
-$\mathbf{F}_{i}^{\text{out}}=\text{GELU}\left(\mathcal{F}_{\text{S}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus\mathcal{F}_{\text{T}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\right)\cdot\mathbf{W}_{\text{up}}.$ (1)
+$\vspace{-3pt}\mathbf{F}_{i}^{\text{out}}=\text{GELU}\left(\mathcal{F}_{\text{S}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus\mathcal{F}_{\text{T}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\right)\cdot\mathbf{W}_{\text{up}}.$ (1)
 Herein, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{C\times C^{\prime}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{C^{\prime}\times C}$ are the transformation matrices of two linear layers for downsampling and upsampling, respectively.
 $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ denote the disentangled feature adaptation by the spatial and temporal pathways respectively.
-Below we discuss two feasible modeling mechanisms for $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$, which leverages 3D convolution and our proposed Anisotropic Deformable Spatio-Temporal Attention, respectively.
+Below we discuss two feasible modeling mechanisms for $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$, which leverages 3D convolution and our proposed aDSTA module, respectively.
 Modeling with 3D Convolution.
@@
 The resulting version is called DST-Adapter.
-While it has been validated that leveraging 3D convolutions with different kernel shapes can learn the spatial and temporal features separately [8, 38], An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain.
-Constructing deep convolutional network by stacking convolutional layers can enlarge the receptive field whilst increasing the model size and incurring heavy computational burden, which is inconsistent with the lightweight design principle for adapters.
-Modeling with Anisotropic Deformable Spatio-Temporal Attention.
+An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain, especially with shallow layers for lightweight design.
+Modeling with aDSTA.
 To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise anisotropic Deformable Spatio-Temporal Attention (aDSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency.
 It adapts the deformable attention [40] from 2D image space to 3D spatio-temporal space.
-As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling density in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions.
+As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling densities in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions.
 As a result, aDSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
@@
 Thus, the feature map $\mathbf{F}^{\prime}_{i}\in\mathbb{R}^{T\times H\times W\times C^{\prime}}$ fused with the position information at the $i$-th stage is obtained by:
-$\mathbf{F}^{\prime}_{i}=\mathcal{F}_{\text{DPE}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}}).$ (2)
+$\vspace{-3pt}\mathbf{F}^{\prime}_{i}=\mathcal{F}_{\text{DPE}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}}).$ (2)
 Our aDSTA learns the offsets in 3D space for the reference points by a 3D convolutional network consisting of a 3D depth-wise convolution layer and a $1\times 1\times 1$ 3D convolution layer as well as a GELU in between.
@@
 For instance, the features for the shifted point $\mathbf{p}$ located at $(p_{t},p_{h},p_{w})$ is calculated via the trilinear interpolation $\mathcal{F}_{\text{tri-int}}$ by:
-$\mathcal{F}_{\text{tri-int}}(\mathbf{p})=\sum_{\mathbf{r}}g(p_{t},r_{t})\cdot g(p_{h},r_{h})\cdot g(p_{w},r_{w})\cdot\mathbf{F}^{\prime}_{i}(\mathbf{r}),$ (3)
+$\vspace{-5pt}\mathcal{F}_{\text{tri-int}}(\mathbf{p})=\sum_{\mathbf{r}}g(p_{t},r_{t})\cdot g(p_{h},r_{h})\cdot g(p_{w},r_{w})\cdot\mathbf{F}^{\prime}_{i}(\mathbf{r}),$ (3)
 where $g(a,b)=\max(0,1-|a-b|)$ defines an interpolating cubic volume space centered at $a$ and $\mathbf{r}=(r_{t},r_{h},r_{w})$ indexes all tokens in the whole spatio-temporal 3D space.
@@
 For instance, a token in $\mathbf{F}^{\prime}_{i}$ located at $(u_{t},u_{h},u_{w})$ serves as a query and the output $\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}$ of aDSTA is calculated by:
-$\begin{split}&\mathbf{q}=\mathbf{F}^{\prime}_{i,(u_{t},u_{h},u_{w})}\cdot\mathbf{W}_{q},\mathbf{K}=\mathbf{P}\cdot\mathbf{W}_{k},\mathbf{V}=\mathbf{P}\cdot\mathbf{W}_{v},\\ &\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}=\text{softmax}(\mathbf{q}\mathbf{K}/\sqrt{C^{\prime}})\mathbf{V}\cdot\mathbf{W}_{o},\end{split}$ (4)
+$\vspace{-3pt}\begin{split}&\mathbf{q}=\mathbf{F}^{\prime}_{i,(u_{t},u_{h},u_{w})}\cdot\mathbf{W}_{q},\mathbf{K}=\mathbf{P}\cdot\mathbf{W}_{k},\mathbf{V}=\mathbf{P}\cdot\mathbf{W}_{v},\\ &\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}=\text{softmax}(\mathbf{q}\mathbf{K}/\sqrt{C^{\prime}})\mathbf{V}\cdot\mathbf{W}_{o},\end{split}$ (4)
 where $\mathbf{P}\in\mathbb{R}^{M\times C^{\prime}}$ is feature matrix of $M$ shifted points and $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ are projection matrices for the query, key, value and output respectively.
-Anisotropic Sampling Density.
+Anisotropic sampling densities.
 The learned offsets for reference points are restricted in a limited range for two reasons.
@@
 Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region.
-As a result, more sampling of reference points typically leads to more powerful and fine-grained representations by the shifted points.
-The deformable attention samples reference points uniformly in the 2D feature space with the same sampling density along two spatial dimensions.
-In contrast, we configure anisotropic sampling density in the spatial and temporal domains.
+As a result, denser sampling of reference points typically leads to finer-grained representations.
+The classical deformable attention samples reference points uniformly in 2D feature space with the same sampling densities along two spatial dimensions.
+In contrast, we configure anisotropic sampling densities in the spatial and temporal domains.
 Consequently, our D2ST-Adapter can leverage aDSTA to model both spatial feature adaptation $\mathcal{F}_{\text{S}}$ in the spatial pathway and temporal feature adaptation $\mathcal{F}_{\text{T}}$ in the temporal pathway.
-Specifically, we define the ‘Sampling kernel’ for aDSTA as:
-Definition 1
-Sampling kernel: formulated as the sampling densities of reference points along the dimensions of time, height and width, respectively, denoted as $<n_{t},n_{s},n_{s}>$, s.t.
-$M=n_{t}\times n_{s}\times n_{s}$.
-$M$ is the total number of sampled reference points.
-Intuitively, an aDSTA module whose sampling kernel has large $n_{t}$ and small $n_{s}$ are more capable of learning the temporal features than learning the spatial features.
+We denote the sampling densities of reference points of aDSTA along the dimensions of time, height and width as $<n_{t},n_{s},n_{s}>$, thereby the total number of sampled reference points is $n_{t}\times n_{s}\times n_{s}$.
+Intuitively, an aDSTA module which samples large $n_{t}$ and small $n_{s}$ is more capable of learning the temporal features than learning the spatial features.
 On the other hand, sampling more reference points in the spatial domain than the temporal domain ($n_{s}>n_{t}$) generally makes aDSTA focus on learning the spatial features.
-Thus, we can tailor aDSTA by configuring the sampling kernel to model $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ correspondingly.
-In particular, we tailor two versions of aDSTA:
-  • aDSTA-S for modeling $\mathcal{F}_{\text{S}}$ in the spatial pathway, whose sampling kernel has relatively large $n_{s}$ and small $n_{t}$ ($n_{s}>n_{t}$).
+Thus, we can tailor aDSTA by configuring the sampling densities to model $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ correspondingly:
+  • aDSTA-S for modeling $\mathcal{F}_{\text{S}}$ in the spatial pathway, which is sampled with larger $n_{s}$ and smaller $n_{t}$ ($n_{s}>n_{t}$).
   • aDSTA-T for modeling $\mathcal{F}_{\text{T}}$ in the temporal pathway, which samples denser reference points in temporal domain than in spatial domain ($n_{t}>n_{s}$).
-In practice, the values of $n_{t}$ and $n_{s}$ are tuned as hyper-parameters.
-Such modeling is analogous to classical property of 3D convolutional network that configuring the shape of the convolutional kernel can steer the convolutional operation to focus on learning features in different domain.
+The values of $n_{t}$ and $n_{s}$ are tuned as hyper-parameters.
 3.3 End-to-End Adapter Tuning
 The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy.
-It first calculates the frame-wise $L_{2}$ distance matrix between the input query video and each class prototype derived by simply averaging the corresponding support samples.
+Following ProtoNet [26], it first calculates the frame-wise $L_{2}$ distance matrix between the query video and each class prototype derived by averaging the corresponding support samples.
 Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction.
 Formally, the similarity between query $q$ and class prototype $c$ can be expressed as:
-$s(\mathbf{F}_{q},\mathbf{F}_{c})=\mathbf{M}([\mathbf{F}_{q}^{1},\cdots,\mathbf{F}_{q}^{T}],[\mathbf{F}_{c}^{1},\cdots,\mathbf{F}_{c}^{T}]),$ (5)
-where $\mathbf{F}_{q}^{i}$ and $\mathbf{F}_{c}^{j}$ denote the features of the $i$-th frame in query $q$ and the $j$-th frame in class prototype $c$ while $\mathbf{M}$ is the matching metric such as OTAM [2] and Bi-MHM [44].
-The whole model can be optimized in an end-to-end manner using the cross-entropy loss.
+$\vspace{-3pt}s(\mathbf{F}_{q},\mathbf{F}_{c})=\mathbf{M}([\mathbf{F}_{q}^{1},\cdots,\mathbf{F}_{q}^{T}],[\mathbf{F}_{c}^{1},\cdots,\mathbf{F}_{c}^{T}]),$ (5)
+where $\mathbf{F}_{q}^{i}$ and $\mathbf{F}_{c}^{j}$ denote the features of the $i$-th frame in query $q$ and the $j$-th frame in class prototype $c$, and $\mathbf{M}$ is the matching metric such as Bi-MHM [33].
+During training, we take the similarities for each class as logits and optimize the model in an end-to-end manner using the same cross-entropy loss as in [3, 23, 33].
+For inference, we classify the query as the support class closest to it.
 4 Experiments
@@
 We conduct experiments on five standard few-shot action recognition benchmarks, including SSv2-Full [8], SSv2-Small [8], Kinetics [14], HMDB51 [15], and UCF101 [27].
-The temporal features are crucial to the action recognition in SSv2-Full and SSv2-Small while the actions in other datasets relies more on the spatial features than the dynamic features.
-Following the typical data split [59, 2, 43], For SSv2-Full, SSv2-Small and Kinetics, we select 64/24/12 classes from the datasets as the training/validation/test set, respectively.
+Following the typical data split [51, 3, 34], for SSv2-Full, SSv2-Small and Kinetics, we select 64/12/24 classes from the datasets as the training/validation/test set, respectively.
 As for HMDB51 and UCF101, we adopt the same data split as [45, 34].
-Pre-trained backbone.
-Most prior arts employs the ResNet-50 [13] pre-trained on ImageNet [5] as backbone to extract frame features.
-Recently, CLIP-FSAR [42] conduct experiments on CLIP-ViT-B [7] and get remarkable performance.
-To verify the effectiveness of our D2ST-Adapter, we instantiate our method with both pre-trained backbones.
+[…]
+Method Pre-training Fine-tuning SSv2-Full SSv2-Small Kinetics HMDB51 UCF101
+1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
+CMN [51] INet-RN50 Full 34.4 43.8 33.4 46.5 57.3 76.0 -- -- -- --
+OTAM [3] INet-RN50 Full 42.8 52.3 36.4 48.0 73.0 85.8 54.5 68.0 79.9 88.9
+AMeFu-Net [7] INet-RN50 Full -- -- -- -- 74.1 86.8 60.2 75.5 85.1 95.5
+ITANet [46] INet-RN50 Full 49.2 62.3 39.8 53.7 73.6 84.3 -- -- -- --
+TRX [23] INet-RN50 Full 42.0 64.6 36.0 59.1 63.6 85.9 53.1 75.6 78.2 96.1
+TA2[18] INet-RN50 Full 47.6 61.0 -- -- 72.8 85.8 59.7 73.9 81.9 95.1
+MTFAN [38] INet-RN50 Full 45.7 60.4 -- -- 74.6 87.4 59.0 74.6 84.8 95.1
+STRM [28] INet-RN50 Full 43.1 68.1 37.1 55.3 62.9 86.7 52.3 77.3 80.5 96.9
+HyRSM [33] INet-RN50 Full 54.3 69.0 40.6 56.1 73.7 86.1 60.3 76.0 83.9 94.7
+Nguyen et al. [20] INet-RN50 Full 43.8 61.1 -- -- 74.3 87.4 59.6 76.9 84.9 95.9
+Huang et al. [12] INet-RN50 Full 49.3 66.7 38.9 61.6 73.3 86.4 60.1 77.0 71.4 91.0
+HCL [48] INet-RN50 Full 47.3 64.9 38.7 55.4 73.7 85.8 59.1 76.3 82.5 93.9
+SA-CT [47] INet-RN50 Full 48.9 69.1 -- -- 71.9 87.1 60.4 78.3 85.4 96.4
+SloshNet [42] INet-RN50 Full 46.5 68.3 -- -- 70.4 87.0 59.4 77.5 86.0 97.1
+GgHM [43] INet-RN50 Full 54.5 69.2 -- -- 74.9 87.4 61.2 76.9 85.2 96.3
+MoLo [34] INet-RN50 Full 56.6 70.6 42.7 56.4 74.0 85.6 60.8 77.4 86.0 95.5
+Zheng et al. [49] INet-RN50 Full 55.4 71.4 41.3 57.5 75.6 87.6 61.5 76.0 87.8 96.0
+ST-Adapter [21] INet-RN50 PEFT 52.2 68.7 41.9 55.7 73.0 85.1 60.3 74.7 84.6 94.5
+D2ST-Adapter INet-RN50 PEFT 57.0 73.6 45.8 60.9 75.8 87.7 61.6 76.6 86.9 95.6
+AIM [44] CLIP-ViT-B PEFT 63.7 79.2 52.8 67.5 88.4 95.3 74.2 86.9 95.4 98.5
+DUALPATH [22] CLIP-ViT-B PEFT 64.5 79.8 53.5 68.1 88.8 95.4 74.9 87.5 95.7 98.7
+ST-Adapter [21] CLIP-ViT-B PEFT 64.2 79.5 53.1 68.0 88.5 95.1 74.1 87.3 95.9 98.9
+D2ST-Adapter CLIP-ViT-B PEFT 66.7 81.9 55.0 69.3 89.3 95.5 77.1 88.2 96.4 99.1
+CLIP-FSAR [35] CLIP-ViT-B Full 62.1 72.1 54.6 61.8 94.8 95.4 77.1 87.7 97.0 99.1
+EMP-Net [37] CLIP-ViT-B PEFT 63.1 73.0 57.1 65.7 89.1 93.5 76.8 85.8 94.3 98.2
+MA-FSAR [41] CLIP-ViT-B PEFT 63.3 72.3 59.1 64.5 95.7 96.0 83.4 87.9 97.2 99.2
+Task-Adapter [2] CLIP-ViT-B PEFT 71.3 74.2 60.2 70.2 95.0 96.8 83.6 88.8 98.0 99.4
+D2ST-Adapter-MM CLIP-ViT-B PEFT 74.1 83.6 63.6 71.9 96.2 97.2 84.3 89.0 98.3 99.5
 Implementation details.
-We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B) as shown in Figure 2 (a).
-We tune the sampling kernel for aDSTA-S and aDSTA-T for each backbone and the details are provided in the supplementary material.
-For a fair comparison with previous methods [2, 44, 43], we uniformly sample 8 frames (i.e., $T=8$) as the input of a video.
-We employ several standard techniques for data augmentation such as random crop and color jitter in the training stage.
-Bi-MHM [44] is used as the matching metric in comparison with other methods.
-Besides, OTAM [2] and TRX [28] are additionally used in the ablation study.
-We train the model episodically with Adam [18] optimizer.
-Following the common practice [44, 43], in the test stage, we randomly select 10,000 episodes from the test set and report the average accuracy.
-[…]
-Method Backbone Adaptation SSv2-Full SSv2-Small
-paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot
-CMN [59] ResNet-50 Full fine-tuning 34.4 -- -- -- 43.8 33.4 42.5 46.5
-CMN-J [60] ResNet-50 Full fine-tuning -- -- -- -- -- 36.2 44.6 48.8
-OTAM [2] ResNet-50 Full fine-tuning 42.8 49.1 51.5 52.0 52.3 36.4 45.9 48.0
-ITANet [56] ResNet-50 Full fine-tuning 49.2 55.5 59.1 61.0 62.3 39.8 49.4 53.7
-TRX [28] ResNet-50 Full fine-tuning 42.0 53.1 57.6 61.1 64.6 36.0 51.9 59.1
-TA2[22] ResNet-50 Full fine-tuning 47.6 -- -- -- 61.0 -- -- --
-MTFAN [46] ResNet-50 Full fine-tuning 45.7 -- -- -- 60.4 -- -- --
-STRM [36] ResNet-50 Full fine-tuning 43.1 53.3 59.1 61.7 68.1 37.1 49.2 55.3
-HyRSM [44] ResNet-50 Full fine-tuning 54.3 62.2 65.1 67.9 69.0 40.6 52.3 56.1
-Nguyen et al. [24] ResNet-50 Full fine-tuning 43.8 -- -- -- 61.1 -- -- --
-Huang et al. [15] ResNet-50 Full fine-tuning 49.3 -- -- -- 66.7 38.9 -- 61.6
-HCL [57] ResNet-50 Full fine-tuning 47.3 54.5 59.0 62.4 64.9 38.7 49.1 55.4
-SloshNet [49] ResNet-50 Full fine-tuning 46.5 -- -- -- 68.3 -- -- --
-GgHM [50] ResNet-50 Full fine-tuning 54.5 -- -- -- 69.2 -- -- --
-MoLo [43] ResNet-50 Full fine-tuning 56.6 62.3 67.0 68.5 70.6 42.7 52.9 56.4
-ST-Adapter [25] ResNet-50 Adapter tuning 52.2 59.9 64.1 67.1 68.7 41.9 51.5 55.7
-D2ST-Adapter (Ours) ResNet-50 Adapter tuning 57.0 65.2 69.5 71.4 73.6 45.8 56.6 60.9
-CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 61.9 64.9 68.1 70.9 72.1 54.5 58.6 61.8
-AIM [51] CLIP-ViT-B Adapter tuning 63.7 72.0 75.2 77.3 79.2 52.8 64.0 67.5
-DUALPATH [26] CLIP-ViT-B Adapter tuning 64.5 72.6 76.0 78.0 79.8 53.5 64.4 68.1
-ST-Adapter [25] CLIP-ViT-B Adapter tuning 64.2 72.4 76.0 77.4 79.5 53.1 64.2 68.0
-D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 66.7 75.3 78.3 80.1 81.9 55.0 65.8 69.3
+To evaluate D2ST-Adapter, we instantiate our method with both ResNet-50 [10] and CLIP-ViT-B/16 [5].
+We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B/16), as shown in Figure 2 (a).
+For a fair comparison with previous methods [33, 34], we uniformly sample 8 frames (i.e.
+$T=8$) as the input of a video.
+Bi-MHM [33] is used as the matching metric in comparison with state-of-the-art methods.
+Besides, OTAM [3] and TRX [23] are additionally employed in the ablation study.
+We report the average classification accuracy of 10,000 episodes randomly selected from the test set.
+Other implementation details follow HyRSM [33], which are provided in the supplementary material.
+4.2 Comparison with State-of-the-Art Methods
+We compare our method with other methods on two types of datasets: temporal-related benchmarks including SSv2-Full and SSv2-Small datasets, and spatial-related benchmarks inlcuding Kinetics, HMDB51, and UCF101 datasets.
+Benchmarks sensitive to temporal features.
+Table 1 shows the comparative results of different methods on two temporal-related datasets, i.e.
+SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition.
+Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin in most few-shot settings.
+These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition.
+In particular, our D2ST-Adapter substantially outperforms all adapter tuning methods for video data including AIM, DUALPATH, and ST-Adapter, highlighting the advantages of its core design, namely built-in disentangled adaptation of spatial and temporal features.
+Benchmarks relying on spatial features.
+Experimental results on the spatial-related benchmarks including Kinetics, HMDB51, and UCF101, are also presented in Table 1.
+Our method still performs best in most settings, which manifest the effectiveness and robustness of our method in these scenarios.
+Besides, the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets.
+It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
+Table 2: Comparison with previous adapter-based methods on traditional full-shot action recognition task with CLIP-ViT-B/16 backbone.
+The testing views (# frames per clip $\times$ # temporal clip $\times$ # spatial crop) are set as 8$\times$1$\times$3 for K400 and 8$\times$3$\times$1 for SSv2.
+Method K400 SSv2
+AIM [44] 83.9 66.4
+DUALPATH [22] 84.1 67.3
+ST-Adapter [21] 82.0 67.1
+D2ST-Adapter (Ours) 84.5 68.0
+Comparison with multimodal methods.
+To have a fair comparison with the multimodal methods for few-shot action recognition, we extend our model to multi-modality (D2ST-Adapter-MM).
+This extension draws on the multimodal framework of CLIP-FSAR [35] to align with other methods.
+Table 1 shows that our method outperforms all other multimodal methods, especially on SSv2-Full and SSv2-Small datasets.
+These findings consistently highlight the effectiveness and robustness of our D2ST-Adapter.
 Figure 3: Illustration of actions our method excels at (1-4) and performs poorly at (5-8), respectively. The performance is evaluated on SSv2-Small benchmark in 1-shot setting.
+Comparison on traditional (full-shot) action recognition task.
+Our model is particularly effective in the few-shot scenario while it can also be applied to traditional action recognition tasks using full training samples.
+To obtain a more comprehensive assessment of our model, we conduct experiments on two traditional action recognition benchmarks, Kinetics-400 (K400) [14] and Something-Something-v2 (SSv2) [8], using the same settings as ST-Adapter [21].
+Table 2 shows that our method still outperforms these adapter-based methods on both benchmarks.
 Figure 4: Performance and efficiency of different methods with ResNet-50 backbone on SSv2-Small benchmark in 1-shot setting. Circle size indicates the memory usage.
-4.2 Comparison with State-of-the-Art Methods
-Benchmarks sensitive to temporal features.
-Table 1 shows the comparative results of different methods on SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition.
-Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin with both backbones in all few-shot settings except ‘5-shot’ on SSv2-Small.
-These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition.
-In particular, our D2ST-Adapter outperforms all existing adapter tuning methods for video data, including AIM, DUALPATH, and ST-Adapter, substantially on all settings, which reveals the benefit of two core technical designs of D2ST-Adapter, namely the dual-pathway architecture for disentangled learning of spatial and temporal features and anisotropic Deformable Spatio-Temporal Attention (aDSTA) for learning features in a global view.
-To gain more insight into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter.
-Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing poorly on simple actions that can be recognized primarily based on spatial features.
-Such results are consistent with the previous results that our model excels at recognizing actions involving complex temporal dynamics.
-[…]
-Method Backbone Adaptation Kinetics HMDB51 UCF101
-paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot 1-shot 3-shot 5-shot
-CMN [59] ResNet-50 Full fine-tuning 57.3 67.5 72.5 74.7 76.0 -- -- -- -- -- --
-CMN-J [60] ResNet-50 Full fine-tuning 60.5 70.0 75.6 77.3 78.9 -- -- -- -- -- --
-OTAM [2] ResNet-50 Full fine-tuning 73.0 75.9 78.7 81.9 85.8 54.5 65.7 68.0 79.9 87.0 88.9
-ITANet [56] ResNet-50 Full fine-tuning 73.6 -- -- -- 84.3 -- -- -- -- -- --
-TRX [28] ResNet-50 Full fine-tuning 63.6 76.2 81.8 83.4 85.9 53.1 66.8 75.6 78.2 92.4 96.1
-TA2[22] ResNet-50 Full fine-tuning 72.8 -- -- -- 85.8 59.7 -- 73.9 81.9 -- 95.1
-MTFAN [46] ResNet-50 Full fine-tuning 74.6 -- -- -- 87.4 59.0 -- 74.6 84.8 -- 95.1
-STRM [36] ResNet-50 Full fine-tuning 62.9 76.4 81.1 83.8 86.7 52.3 67.4 77.3 80.5 92.7 96.9
-HyRSM [44] ResNet-50 Full fine-tuning 73.7 80.0 83.5 84.6 86.1 60.3 71.7 76.0 83.9 93.0 94.7
-Nguyen et al. [24] ResNet-50 Full fine-tuning 74.3 -- -- -- 87.4 59.6 -- 76.9 84.9 -- 95.9
-Huang et al. [15] ResNet-50 Full fine-tuning 73.3 -- -- -- 86.4 60.1 -- 77.0 71.4 -- 91.0
-HCL [57] ResNet-50 Full fine-tuning 73.7 79.1 82.4 84.0 85.8 59.1 71.2 76.3 82.5 91.0 93.9
-SloshNet [49] ResNet-50 Full fine-tuning 70.4 -- -- -- 87.0 59.4 -- 77.5 86.0 -- 97.1
-GgHM [50] ResNet-50 Full fine-tuning 74.9 -- -- -- 87.4 61.2 -- 76.9 85.2 -- 96.3
-MoLo [43] ResNet-50 Full fine-tuning 74.0 80.4 83.7 84.7 85.6 60.8 72.0 77.4 86.0 93.5 95.5
-ST-Adapter [25] ResNet-50 Adapter tuning 73.0 79.9 82.8 84.6 85.1 60.3 71.4 74.7 84.6 92.9 94.5
-D2ST-Adapter (Ours) ResNet-50 Adapter tuning 75.0 81.4 84.3 85.8 87.0 61.6 73.0 76.6 86.9 94.4 95.6
-CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 89.7 92.9 94.2 94.8 95.0 75.8 84.1 87.7 96.6 98.4 99.0
-AIM [51] CLIP-ViT-B Adapter tuning 88.4 92.7 94.2 94.6 95.3 74.2 83.8 86.9 95.4 98.2 98.5
-DUALPATH [26] CLIP-ViT-B Adapter tuning 88.8 92.8 94.3 94.7 95.4 74.9 84.6 87.5 95.7 98.4 98.7
-ST-Adapter [25] CLIP-ViT-B Adapter tuning 88.5 92.6 94.0 94.6 95.1 74.1 84.3 87.3 95.9 98.5 98.9
-D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 89.3 93.1 94.4 95.2 95.5 77.1 86.4 88.2 96.4 98.7 99.1
-Benchmarks relying on spatial features.
-Table 2 presents the experimental results on the benchmarks relying more on spatial features, including Kinetics, HMDB51 and UCF101.
-Our method still performs best on most of the settings, which manifest the effectiveness and robustness of our method in these scenarios.
-We also observe that the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets.
-It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
+Case Study of the Pros and Cons of D2ST-Adapter.
+To gain more insights into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels at, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter.
+Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing on par with other methods on simple actions that can be recognized primarily based on spatial features.
+Such results are consistent with the conclusion that our model excels at recognizing actions involving complex temporal dynamics.
+Table 3: Comparison with other adapter-based methods using CLIP-ViT-B/16 as backbone on efficiency. Accuracy on SSv2-Full benchmark in 1-shot setting is also provided.
+Method Tunable Memory GFLOPs Latency Acc.
+Params Usage
+AIM [44] 14.3 M 19.0 GB 208 14.1 ms 63.7
+DUALPATH [22] 14.3 M 16.7 GB 171 11.6 ms 64.5
+ST-Adapter [21] 14.3 M 16.1 GB 163 12.6 ms 64.2
+D2ST-Adapter 7.3 M 17.3 GB 150 18.2 ms 66.7
 Comparison of efficiency.
@@
 It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation.
-Meanwhile, our method outperforms these full fine-tuning based methods substantially while enjoying much fewer tunable parameters, which validates the superiority of our method.
-4.3 Ablation study
-We conduct ablation study on SSv2-Full and Kinetics datasets, using ResNet-50 as the pre-trained backbone and Bi-MHM [44] as the matching metric.
-Comparison between different Adapters.
-We first compare the performance of different Adapters in Table 3.
-As shown in Figure 1, Vanilla-Adapter [14] only contains the downsampling and upsampling layers with a GELU nonlinearity in between.
-As described in Section 3.2, DST-Adapter is the convolutional version of our D2ST-Adapter which follows the dual-pathway structure but models spatial and temporal pathways with 3D convolution.
-The performance of full fine-tuning is also provided for reference.
-We make following observations.
-  • Effect of learning temporal features by adapters.
-ST-Adapter outperforms Vanilla-Adapter distinctly, which indicates that using 3D convolution to learn joint spatio-temporal features facilitates the feature adaptation.
-  • Effect of disentangled encoding of the spatial and temporal features.
-The performance improvement from ST-Adapter to DST-Adapter demonstrates the effectiveness of dual-pathway adapter architecture, which allows for encoding the spatial and temporal features in a disentangled manner.
-  • Effect of the proposed anisotropic Deformable Spatial-Temporal Attention (aDSTA).
-Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on the SSv2-Full dataset, which is more challenging due to the necessity of encoding temporal features.
-Table 3: Comparison between different adapters as well as full fing-tuning.
+Meanwhile, our method outperforms these full fine-tuning based methods substantially with much lower parameter-finetuning overhead, which validates the superiority of our method.
+We further compare our model with other adapter-based methods on efficiency in Table 3.
+Our D2ST-Adapter performs best while maintaining relatively high efficiency due to the lightweight design of the proposed aDSTA module and the smaller number of inserted adapters (only one per block in our method vs.
+two or four per block in other adapter-based methods [21, 44, 22]).
+Figure 5: Visualization of important shifted reference points w.r.t all queries (first sample), and important shifted reference points w.r.t a given query, marked with a red star (second sample).
+More visualization results are provided in the supplementary material.
+Table 4: Comparison between different adapters.
 Method SSv2-Full Kinetics
@@
 ST-Adapter 52.2 68.7 73.0 85.1
-DST-Adapter (Ours) 53.7 69.6 73.3 85.5
-D2ST-Adapter (Ours) 57.0 73.6 75.0 87.0
-Table 4: Ablation study on the configuration of sampling kernel of aDSTA.
-Configuration SSv2-Full Kinetics
+DST-Adapter (Ours) 53.7 69.6 74.4 86.4
+D2ST-Adapter (Ours) 57.0 73.6 75.8 87.7
+Table 5: Effect of the anisotropic sampling of aDSTA.
+Method Configurations SSv2-Full Kinetics
 1-shot 5-shot 1-shot 5-shot
+aDSTA-S 55.2 71.9 74.1 86.0
+Single branch aDSTA-T 56.3 72.5 73.8 85.8
 aDSTA-Uniform 55.8 72.2 73.9 86.0
-aDSTA-S & aDSTA-T 57.0 73.6 75.0 87.0
+Dual branches aDSTA-S & aDSTA-T 57.0 73.6 75.8 87.7
+4.3 Ablation study
+We conduct ablation study on SSv2-Full and Kinetics datasets, using ResNet-50 as the pre-trained backbone and Bi-MHM [33] as the matching metric.
+Comparison between different Adapters.
+We first compare the performance of different Adapters in Table 4.
+Vanilla-Adapter [11] only contains the downsampling and upsampling layers with a GELU nonlinearity in between.
+As described in Section ‘Method’, DST-Adapter is the convolutional version of our D2ST-Adapter which models both the spatial and temporal pathways with 3D convolution.
+The performance of full fine-tuning is also provided for reference.
+We make following observations.
+  • Effect of learning temporal features by adapters.
+ST-Adapter outperforms Vanilla-Adapter distinctly, which indicates that using 3D convolution to learn joint spatio-temporal features facilitates the feature adaptation.
+  • Effect of disentangled encoding of the spatial and temporal features is demonstrated by the performance improvement from ST-Adapter to DST-Adapter benefited from the dual-pathway adapter architecture.
+  • Effect of the proposed aDSTA.
+Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on SSv2-Full which demands effective encoding of temporal features.
+Effect of anisotropic sampling of aDSTA.
+To investigate the effect of proposed anisotropic sampling scheme of aDSTA, we compare the performance of our D2ST-Adapter using four different configurations of the aDSTA: 1) only ‘aDSTA-S’ modeling the spatial pathway is used; 2) only ‘aDSTA-T’ for the temporal pathway is preserved; 3) ‘aDSTA-Uniform’ performs sampling with uniform density in both the spatial and temporal domains; and 4) ‘aDSTA-S & aDSTA-T’ corresponds the fully functional D2ST-Adapter.
+Table 5 shows that the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling densities for the spatial and temporal pathways.
 Table 6: Evaluation of generalization of our D2ST-Adapter across different matching metrics on SSv2-Full and Kinetics. The state-of-the-art methods for each metric are involved into comparison.
@@
 1-shot 5-shot 1-shot 5-shot
-RFPL [47] 47.0 61.0 74.6 86.8
-OTAM [2] MoLo [43] 55.0 69.6 73.8 85.1
+RFPL 47.0 61.0 74.6 86.8
+OTAM MoLo 55.0 69.6 73.8 85.1
 D2ST-Adapter 56.0 72.8 74.7 86.8
-RFPL [47] 44.6 64.6 66.2 87.3
-TRX [28] MoLo [43] 45.6 66.1 64.8 86.3
+RFPL 44.6 64.6 66.2 87.3
+TRX MoLo 45.6 66.1 64.8 86.3
 D2ST-Adapter 47.7 68.6 66.4 87.6
-HyRSM [44] 54.3 69.0 73.7 86.1
-Bi-MHM [44] MoLo [43] 56.6 70.6 74.0 85.6
-D2ST-Adapter 57.0 73.6 75.0 87.0
-Effect of configuration of the sampling kernel of aDSTA.
-To investigate the effect of configuring specialized sampling kernels for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway respectively, we additionally configure an aDSTA-Uniform module, which performs sampling with uniform density in both the spatial and temporal domains.
-Then we compare the performance of our D2ST-Adapter using two different settings, including 1) ‘aDSTA-Uniform’, in which aDSTA-Uniform is used to model both the spatial and temporal pathways and 2) ‘aDSTA-S & aDSTA-T’, in which D2ST-Adapter uses aDSTA-S and aDSTA-T to model the spatial and temporal pathways respectively.
-As shown in Table 4, the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling kernels for the spatial and temporal pathways.
+HyRSM 54.3 69.0 73.7 86.1
+Bi-MHM MoLo 56.6 70.6 74.0 85.6
+D2ST-Adapter 57.0 73.6 75.8 87.7
+Generalization Across Different Matching Metrics.
+To evaluate the generalization of our D2ST-Adapter across different matching metrics, we conduct the experiments using three classical matching metrics on SSv2-Full and Kinetics, including OTAM [3], TRX [23], and Bi-MHM [33], and compare our model with the state-of-the-art methods corresponding to each metric.
+Table 6 presents the results, which show that our D2ST-Adapter consistently outperforms other methods, especially on the challenging SSv2-Full benchmark, validating the well generalization of the proposed D2ST-Adapter across different matching metrics.
+Visualization.
+We visualize the shifted reference points learned by our D2ST-Adapter to investigate the proposed aDSTA qualitatively.
+Following Deformable Attention [40], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance.
+Then we visualize top-100 most important shifted reference points in each pathway, with circle size indicating their importance, as shown in the first sample of Figure 5.
+We also provide visualization of the most relevant reference points w.r.t a certain salient region (token) in the second sample of Figure 5.
+To be specific, we manually select a salient region (indicated by a red star marker) as the query token and visualize top-50 relevant shifted reference points to the query in terms of attention weights.
+The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways.
+Another observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamics.
+This is consistent with the anisotropic sampling of aDSTA for different pathways.
+5 Concluding Remarks
+We present D2ST-Adapter, which is a novel adapter tuning method for few-shot action recognition.
+It is designed in a dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features.
+Moreover, we design the aDSTA module, which enables D2ST-Adapter to encode features in a global view while maintaining lightweight design.
+Extensive experiments demonstrate the superiority of our model over state-of-the-art methods, especially in challenging scenarios involving complex temporal features.
+Supplementary Material
+Appendix A Effect of the Sampling Densities of D2ST-Adapter
+We compare the efficiency and performance of different sampling densities in 1-shot setting in Table 7.
+It demonstrates that sampling denser reference points yields marginal performance gain at the expense of increased memory usage.
+Our model is set to $<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ to strike a good balance between efficiency and performance.
+Appendix B Effect of the Number of Input Frames
+To analyze the impact of the number of input frames, we compare the SSv2-Full 1-shot performance when using different number of frames as input under two sampling densities settings.
+As shown in Table 8, more input frames and denser sampling both yield better performance while incurring more computational cost.
+Appendix C Instantiation on More Backbones
+To investigate the performance of our D2ST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings.
+Table 11 shows that our D2ST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method across different backbones.
+While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to D2ST-Adapter, which manifests the effectiveness of the proposed anisotropic Deformable Spatio-Temporal Attention (aDSTA).
+For larger backbones, we conduct experiments on CLIP-ViT-L/14, which has a total of 303.2 million parameters, significantly larger than CLIP-ViT-B/16, which has a total of 85.8 million parameters.
+The results in Table 12 show that our model can still consistently outperform ST-Adapter and Full Fine-tuning on the larger and stronger backbone.
+Table 7: Effect of the sampling densities of D2ST-Adapter in CLIP-ViT-B.
+Sampling Densities Memory Usage SSv2-Full Kinetics
+(Spatial & Temporal)
+$<$1, 4, 4$>$ &\&& $<$4, 2, 2$>$ 17.0 GB 65.9 89.0
+$<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ 17.3 GB 66.7 89.3
+$<$2, 6, 6$>$ &\&& $<$8, 3, 3$>$ 18.0 GB 66.9 89.2
+$<$2, 8, 8$>$ &\&& $<$8, 4, 4$>$ 18.8 GB 66.9 89.6
+$<$4, 8, 8$>$ &\&& $<$16, 4, 4$>$ 20.3 GB 67.2 89.3
+Table 8: Effect of the number of input frames in SSv2-Full 1-shot setting with CLIP-ViT-B backbone.
+Sampling Densities Sampling Frames
+(Spatial & Temporal) 4 8 16
+$<$1, 4, 4$>$ &\&& $<$4, 2, 2$>$ 63.1 65.9 69.0
+$<$2, 4, 4$>$ &\&& $<$8, 2, 2$>$ 63.6 66.7 69.9
+Table 9: Effect of the inserted position of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset. Skip means adding D2ST-Adapter every other stage.
+Insertion Position Tunable Memory Usage 1-shot 5-shot
+Params (%)
+Early-insertion 4.1% 15.7 GB 48.4 64.6
+Late-insertion 4.1% 15.3 GB 54.2 68.5
+Skip-insertion 4.1% 15.4 GB 53.3 67.9
+Full-insertion 7.9% 17.3 GB 55.0 69.3
+Table 10: Effect of the bottleneck ratio of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset.
+Ratio Tunable Memory Usage 1-shot 5-shot
+Params (%)
+0.0625 1.4% 15.5 GB 53.4 68.1
+0.125 3.2% 16.2 GB 54.2 68.6
+0.25 7.9% 17.3 GB 55.0 69.3
+0.5 20.2% 19.7 GB 54.7 69.4
 Table 11: Performance of different methods using smaller backbones (i.e., ResNet-18 and ResNet-34) on SSv2-Full dataset.
@@
 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 2-shot 3-shot 4-shot 5-shot
-OTAM [2] 39.4 45.0 46.6 47.4 49.0 40.6 45.2 48.0 48.9 49.2
-TRX [28] 29.9 38.2 44.0 48.2 50.3 32.4 41.6 47.7 52.0 53.5
-HyRSM [44] 46.6 54.7 58.7 60.7 61.1 50.0 57.5 61.9 63.3 64.8
-MoLo [43] 50.0 57.2 61.6 63.6 64.6 54.1 61.1 65.9 67.3 67.8
-ST-Adapter [25] 47.3 54.3 58.1 61.3 62.2 48.9 55.6 59.4 62.5 64.1
+OTAM 39.4 45.0 46.6 47.4 49.0 40.6 45.2 48.0 48.9 49.2
+TRX 29.9 38.2 44.0 48.2 50.3 32.4 41.6 47.7 52.0 53.5
+HyRSM 46.6 54.7 58.7 60.7 61.1 50.0 57.5 61.9 63.3 64.8
+MoLo 50.0 57.2 61.6 63.6 64.6 54.1 61.1 65.9 67.3 67.8
+ST-Adapter 47.3 54.3 58.1 61.3 62.2 48.9 55.6 59.4 62.5 64.1
 DST-Adapter (Ours) 50.0 56.9 61.7 63.4 65.3 52.1 59.3 63.4 65.8 67.5
 D2ST-Adapter (Ours) 53.0 60.4 65.0 67.1 68.6 54.4 62.5 66.0 69.1 70.6
-Generalization across different matching metrics.
-To evaluate the generalization of our D2ST-Adapter across different matching metrics, we conduct the experiments using three classical matching metrics on SSv2-Full and Kinetics, including OTAM [2], TRX [28], and Bi-MHM [44], and compare our model with the state-of-the-art methods corresponding to each metric.
-Table 5 presents the results, which show that our D2ST-Adapter consistently outperforms other methods, especially on the challenging SSv2-Full benchmark, validating the well generalization of the proposed D2ST-Adapter across different matching metrics.
-Instantiation on ResNet-18 and ResNet-34.
-To investigate the performance of our D2ST-Adapter on smaller backbones, we instantiate our model with ResNet-18 and ResNet-34 pre-trained on ImageNet, and conduct experiments to compare our model with other methods under the same experimental settings.
-Table 6 show that our D2ST-Adapter achieves the best performance using both backbones in all settings, which demonstrates the robustness of our method.
-While DST-Adapter, the convolutional version of our model, also performs well, it is still inferior to D2ST-Adapter, which manifests the effectiveness of the proposed anisotropic Deformable Spatio-Temporal Attention (aDSTA).
-Figure 5: Visualization of shifted reference points in both pathways for two video samples.
-Circle size indicates the importance for each point calculated by aggregating the attention from all queries.
-4.4 Visualization
-We visualize the shifted reference points learned by our D2ST-Adapter to investigate the effectiveness of the proposed aDSTA qualitatively.
-Following Deformable Attention [48], for each shifted reference point serving as key and value, we accumulate the attention weights assigned by all queries as its relative importance.
-Then we visualize top-100 most important shifted reference points in each pathway for a video sample, as shown in  Figure 5.
-Note that we sample 8 frames for all videos, if a reference point is shifted to a temporal position between two sampled frames, we visualize it on the nearest frame approximately.
-The results show that the shifted reference points learned by our model can always focus on the salient targets that are critical for action recognition in both the spatial and temporal pathways.
-Another interesting observation is that the shifted points in the spatial pathway are distributed densely in some specific frames to capture the spatial appearance features for the targets, while the shifted points in the temporal pathway are distributed uniformly among all frames to capture the temporal dynamic features.
-This is consistent with the design of specialized tailored aDSTA modules for different pathways by configuring the sampling kernels accordingly.
-5 Conclusion
-We present D2ST-Adapter, which is a novel adapter tuning method for few-shot action recognition.
-It is designed in a dual-pathway architecture, which allows for disentangled encoding of spatial and temporal features.
-Moreover, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component of D2ST-Adapter, which enables it to encode features in a global view while maintaining lightweight design.
-Extensive experiments demonstrate that the superiority of our model over state-of-the-art methods, especially in challenging scenarios involving complex temporal features.
-Supplementary Material
-Appendix 0.A Tuning of the Sampling Kernels of aDSTA
-The sampling kernels of aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter can be tuned on a held-out validation set.
-Generally, aDSTA-S should sample denser reference points in the spatial domain while aDSTA-T samples more points in the temporal domain.
-For the instantiation of our model with ResNet, we configure the sampling kernel of aDSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size.
-In contrast, we only need to tune one configuration for sampling kernel when using ViT as the backbone since the feature map always has fixed size in different stages.
-Table 7 shows the tuned configurations of the sampling kernels of aDSTA-S and aDSTA-T with different backbones.
-Besides, we also provide the configurations of sampling kernel for aDSTA-Uniform module which is constructed to validate the effect of configuring the sampling kernel of aDSTA in Table 4.
-Table 7: Tuned sampling kernels (in the form of $(T,H,W)$) for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter.
-Besides, the sampling kernel for aDSTA-Uniform used in Table 4 for ablation study is also provided.
-Backbone Feature Map Sampling Kernel Sampling Kernel Sampling Kernel
+Table 12: Performance of different methods using a larger backbone (CLIP-ViT-L/14). The number of backbone parameters (Params) is also provided as reference.
+Method Backbone (Params) SSv2-Full Kinetics
+1-shot 5-shot 1-shot 5-shot
+Full Fine-tuning CLIP-ViT-L/14 (303.2 M) 55.5 73.0 91.7 96.8
+ST-Adapter CLIP-ViT-L/14 (303.2 M) 67.0 83.8 92.2 96.9
+D2ST-Adapter CLIP-ViT-L/14 (303.2 M) 69.6 85.6 92.8 97.2
+Table 13: Hyper-parameter settings of our method with ResNet-50 backbone.
+Backbone Hyper-parameter SSv2-Full SSv2-Small Kinetics HMDB51 UCF101
+ResNet-50 base learning rate 2e-3 2e-3 2e-3 2e-3 2e-3
+warmup start learning rate 4e-4 4e-4 4e-4 4e-4 4e-4
+training episodes 120000 40000 10000 8000 8000
+warmup episodes 12000 4000 1000 800 800
+weight decay 5e-4 5e-4 5e-4 5e-4 5e-4
+batch size 8 8 8 8 8
+num. adapters per block 1 1 1 1 1
+num. input frames 8 8 8 8 8
+bottleneck ratio 0.25 0.25 0.25 0.25 0.25
+training crop size 224 224 224 224 224
+CLIP-ViT-B base learning rate 1e-3 1e-3 5e-4 1e-3 5e-4
+warmup start learning rate 2e-4 2e-4 1e-4 2e-4 1e-4
+training episodes 120000 40000 3000 3000 2400
+warmup episodes 12000 4000 300 300 240
+weight decay 5e-4 5e-4 5e-4 5e-4 5e-4
+batch size 8 8 8 8 8
+num. adapters per block 1 1 1 1 1
+num. input frames 8 8 8 8 8
+bottleneck ratio 0.25 0.25 0.25 0.25 0.25
+training crop size 224 224 224 224 224
+Table 14: Tuned sampling densities (in the form of $<$T, H, W$>$) for aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter.
+Besides, the sampling densities for aDSTA-Uniform used in Table 5 for ablation study is also provided.
+Backbone Feature Map Sampling Densities Sampling Densities Sampling Densities
 in aDSTA-S in aDSTA-T in aDSTA-Uniform
-ResNet-50 (8, 56, 56) (2, 8, 8) (8, 4, 4) (4, 4, 4)
-ResNet-50 (8, 28, 28) (2, 4, 4) (8, 2, 2) (4, 4, 4)
-ResNet-50 (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
-ResNet-50 (8, 7, 7) (2, 2, 2) (8, 1, 1) (4, 4, 4)
-CLIP-ViT-B (8, 14, 14) (2, 4, 4) (8, 2, 2) (4, 4, 4)
-Appendix 0.B Effect of the Inserted Position of D2ST-Adapter
+ResNet-50 $<$8, 56, 56$>$ $<$2, 8, 8$>$ $<$8, 4, 4$>$ $<$4, 4, 4$>$
+ResNet-50 $<$8, 28, 28$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
+ResNet-50 $<$8, 14, 14$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
+ResNet-50 $<$8, 7, 7$>$ $<$2, 2, 2$>$ $<$8, 1, 1$>$ $<$4, 4, 4$>$
+CLIP-ViT-B $<$8, 14, 14$>$ $<$2, 4, 4$>$ $<$8, 2, 2$>$ $<$4, 4, 4$>$
+Appendix D Effect of the Inserted Position of D2ST-Adapter
 Theoretically, our D2ST-Adapter can be inserted into any position of the backbone flexibly.
@@
 Besides, full-insertion achieves the best performance at the expense of slightly more tunable parameters and memory usage.
-Table 8: Effect of the inserted position of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset. Skip means adding D2ST-Adapter every other stage.
-Insertion Position Tunable Parameters (%) Memory Usage 1-shot 5-shot
-Early-insertion 4.1% 15.5GB 48.4 64.6
-Late-insertion 4.1% 15.2GB 54.2 68.5
-Skip-insertion 4.1% 15.2GB 53.3 67.9
-Full-insertion 7.9% 16.7GB 55.0 69.3
-Table 9: Effect of the bottleneck ratio of D2ST-Adapter in CLIP-ViT-B on SSv2-Small dataset.
-Ratio Tunable Parameters (%) Memory Usage 1-shot 5-shot
-0.0625 1.4% 15.2GB 53.4 68.1
-0.125 3.2% 15.7GB 54.2 68.6
-0.25 7.9% 16.7GB 55.0 69.3
-0.5 20.2% 18.7GB 54.7 69.4
-Appendix 0.C Effect of the Bottleneck Ratio of D2ST-Adapter
+Appendix E Effect of the Bottleneck Ratio of D2ST-Adapter
 The tunable parameter size is mainly determined by the bottleneck ratio of D2ST-Adapter, defined as the ratio of downsampled channel numbers to the initial size.
@@
 As shown in Table 10, larger bottleneck ratios typically yield more performance improvement while introducing more tunable parameters, and we set the bottleneck ratio to 0.25 in all the experiments based on the results.
-Figure 6: Given a selected query within the salient object, top-50 relevant shifted reference points in terms of attention weights in both pathways are visualized for three video samples.
-The red stars denote the selected query tokens while the circles denote relevant shifted reference points.
-Appendix 0.D More Visualizations
-To provide more insight into the anisotropic Deformable Spatio-Temporal Attention (aDSTA), we provide more visualization results in Figure 6.
-To be specific, we manually select a point within the salient object as the query token and visualize top-50 relevant shifted reference points to the query in terms of attention weights.
-The results show that our model is able to capture the salient object through the shifted reference points in both spatial and temporal domains.
+Appendix F More Visualizations
+Consistent with Figure 5 in the paper, we provide more visualization of the shifted reference points in Figures 6 and 7 to evaluate the anisotropic Deformable Spatio-Temporal Attention (aDSTA).
+Note that only the top-100/50 most important points from the aDSTA modules across all inserted D2ST-Adapters are visualized.
+The results show that our model is able to capture the salient objects through the shifted reference points in both spatial and temporal domains.
+Appendix G More Implementation Details
+Matching Metric.
+In the few-shot action recognition task, the classification of a query sample is based on the similarities between it and each support class prototype.
+Typically, a matching metric is first used to temporally align the frames or segments within two videos.
+Then, the overall similarity between the query and each prototype can be calculated by averaging or summing the similarities of all aligned pairs.
+We adopt three classic matching metrics in our experiments, including OTAM [3], TRX [23], and Bi-MHM [33].
+Details of these matching metrics are provided as follows.
+OTAM [3] explicitly leverages the temporal ordering information in videos through tight temporal alignment, which ensures that the alignment results for any two frame pairs do not overlap.
+It extends the Dynamic Time Warping (DTW) algorithm [19] to compute the alignment path of two videos in the temporal dimension, and employs continuous relaxation to make the model differentiable.
+TRX [23] aligns segments rather than individual frames, thus effectively matching actions performed at varying speeds and in different locations across two videos.
+It first exhaustively enumerates all subsequences of two to four frames as potential actions in both videos.
+Then, for each action in the query video, the cross-attention mechanism is employed to compute the corresponding action prototype in the support video based on feature similarities, serving as its match.
+Bi-MHM [33] stands for Bidirectional Mean Hausdorff Metric, which treats the similarity measurement between two videos as a set matching problem.
+This alignment strategy eliminates the constraints of temporal order and focuses solely on the appearance similarities between frames.
+Data Pre-processing.
+We use the same data pre-processing methods as in HyRSM [33].
+A video is uniformly sampled 8 frames as input.
+During training, each frame is first resized to 256$\times$256 and then randomly cropped to 224$\times$224.
+Some basic data augmentation methods are adopted, such as color jitter and horizontal flip.
+Note that for temporal-related datasets, i.e.
+SSv2-Full and SSv2-Small, we do not use horizontal flip, since the recognition of some classes in these two datasets requires distinguishing between left and right, e.g.
+“pulling something from left to right”.
+Tuned hyper-parameters. For ease of the reproduction, we list the tuned values for all the hyper-parameters in our experiments for each dataset with both backbones in Table 13.
+Tuned sampling densities of aDSTA.
+The sampling densities of aDSTA-S in the spatial pathway and aDSTA-T in the temporal pathway of our D2ST-Adapter can be tuned on a held-out validation set.
+Generally, aDSTA-S should sample denser reference points in the spatial domain while aDSTA-T samples more points in the temporal domain.
+For the instantiation of our model with ResNet, we configure the sampling densities of aDSTA in each convolutional stage individually since the feature maps in different convolutional stages may have different size.
+In contrast, we only need to tune one configuration for sampling densities when using ViT as the backbone since the feature map always has fixed size in different stages.
+Table 14 shows the tuned configurations of the sampling densities of aDSTA-S and aDSTA-T with different backbones.
+Besides, we also provide the configurations of sampling densities for aDSTA-Uniform module which is constructed to validate the effect of configuring the sampling densities of aDSTA in Table 5.
+Figure 6: Visualization of top-100 important shifted reference points in both pathways for two video samples.
+Circle size indicates the importance for each point calculated by aggregating the attention weights from all queries.
+Figure 7: Given a selected query within the salient object indicated by a red star marker, top-50 relevant shifted reference points in terms of attention weight are visualized.